### v3

# Generalized Kummer surfaces over finite fields

Abstract.
In this paper, we prove a refinement of the Katsura theorem on finite groups acting on abelian surfaces such that the quotient is birational to a $K3$ surface. As an application, we compute traces of Frobenius on the Neron–Severi groups of supersingular generalized Kummer surfaces over finite fields.
Key words and phrases:
Finite field, abelian variety, Kummer surface
1991 Mathematics Subject Classification:
14K99, 14G05, 14G15
1. Introduction
Throughout this paper $k$ is a perfect field of characteristic $p$, and ${\bar{k}}$ is an algebraic closure of $k$. Let $A$ be an abelian surface over $k$ with an action of a finite group $G$. In this paper, we study only actions such that $G$ preserves the group law on $A$; in other words, there is a homomorphism of the groups $G\to\operatorname{End}(A)^{*}$, where for a ring $R$ we denote by $R^{*}$ the multiplicative group of invertible elements. If a minimal smooth model $X(A,G)$ of the quotient $A/G$ is a $K3$ surface, we say that $X(A,G)$ is a generalized Kummer surface. In [Ry12] we classified zeta functions of Kummer surfaces $X(A,{\mathbb{Z}}/2{\mathbb{Z}})$ over finite fields of characteristic $p>2$. This paper is a step towards such a classification for generalized Kummer surfaces.
For an algebraic variety $X$ over $k$ we denote by $\bar{X}$ the base change of $X$ to ${\bar{k}}$. A $K3$ surface $X$ is called Shioda supersingular if the rank of $\operatorname{NS}(\bar{X})$ is $22$. A Shioda supersingular $K3$ surface is supersingular, and a supersingular Kummer surface is Shioda supersingular [Ar]. For a Shioda supersingular $K3$ surface we have the natural isomorphism for any prime $\ell\neq p$:
()( ∗ ) $\operatorname{NS}(\bar{X})\otimes{\mathbb{Q}}_{\ell}\to H^{2}(\bar{X},{\mathbb{Q}}_{\ell}).$
It follows that, if the base field $k$ is finite, the zeta function of a Shioda supersingular $K3$ surface $X$ is uniquely determined by the Frobenius action on the Neron-Severi group $\operatorname{NS}(\bar{X})$. This observation is a link between the arithmetic and geometry of supersingular $K3$ surfaces.
Since odd cohomology groups of $X$ are trivial, the isomorphism ()(*)( ∗ ) leads to the formula for the number of points on $X$ over $k$:
()( ∗ ∗ ) $|X(k)|=1+q\operatorname{Tr}_{X}+q^{2},$
where $\operatorname{Tr}_{X}$ is the trace of the Frobenius action on the Neron–Severi group of $X$. The main question is: what are the possible values of $\operatorname{Tr}_{X}$, where $X$ runs through all supersingular $K3$ surfaces over ${\mathbb{F}}_{q}$?
Our main motivation comes from the analogy with cubic surfaces in ${\mathbb{P}}^{3}$. For a cubic surface $X$ we have the same formula ()(**)( ∗ ∗ ) for the number of points. In [Serre12] Serre asked which values of $\operatorname{Tr}_{X}$ can arise for smooth cubic surfaces in ${\mathbb{P}}^{3}$ over ${\mathbb{F}}_{q}$. The answer was recently obtained in [BFL]. A more general question on zeta functions of cubic surfaces was treated in [RT16, T20, LT20]. A direct generalization of Serre’s question to quartics in ${\mathbb{P}}^{3}$ seems very difficult, but we can try to generalize the question as follows. Since a quartic in ${\mathbb{P}}^{3}$ with reasonable singularities is a $K3$ surface, we can study traces of Frobenius on Neron–Severi groups of $K3$ surfaces. The isomorphism ()(*)( ∗ ) holds for any smooth cubic surface, but not for any $K3$ surface; therefore, we focus on Shioda supersingular $K3$ surfaces. In this paper, we compute Frobenius traces for generalized Kummer surfaces only, but it is natural to pose the question for supersingular $K3$ surfaces as well.
We heavily use results from the pioneering paper of Katsura on generalized Kummer surfaces over fields of positive characteristics [Ka87]. We reformulate some of his results here in a more convenient way.
Definition 1.1.
We say that the action of a finite group $G$ on an abelian variety $A$ is rigid if the representation of $G$ in $V_{\ell}(A)$ is without fixed points, i.e., for any $g\in G$ of order $r$ the eigenvalues of the action of $g$ on $V_{\ell}(A)$ are primitive roots of unity of degree $r$.
Remark 1.2.
If an action of $G$ on $A$ is rigid, then it is faithful. Moreover, if $g\in G$ is an element of order $2$, then $g$ acts on $A$ as $-1$; thus there is at most one element of order $2$ in $G$.
Theorem 1.3.
[Ka87, Theorem 2.4] Assume that $\mathop{\mathrm{char}}\nolimits k\neq 2$. A relatively minimal model of $A/G$ is a $K3$ surface if and only if $G$ satisfies the following conditions
  • the action is rigid;
  • the action is symplectic;
  • $A/G$ is singular, and all the singular points of $A/G$ are rational double points.
Katsura classified finite groups $G$ such that there exists a finite field $k$ of characteristic $p>5$ and an abelian surface $A$ over $k$ with an action of $G$ such that the quotient is birational to a $K3$ surface.
Theorem 1.4.
[Ka87, Theorem 3.7] Let $k$ be a field of characteristic $p>5$. Let $A$ be an abelian surface over $k$ with an action of a finite group $G$ such that the quotient is birational to a $K3$ surface. If all elements of $G$ fix the group law of $A$, then $G$ is one of the following groups:
  1. (1)
    a cyclic group of order $2,3,4,5,6,8,10$, or $12$;
  2. (2)
    a binary dihedral group $Q_{4n}$ of order $4n$, where $2\leq n\leq 6$;
  3. (3)
    $\operatorname{SL}_{2}({\mathbb{F}}_{3})$, $\operatorname{ESL}_{2}({\mathbb{F}}_{3})$, or $\operatorname{SL}_{2}({\mathbb{F}}_{5})$.
Katsura showed that any group from Theorem 1.4 occurs over some finite field of characteristic $p>5$. As far as author knows, there is no classification of such groups over a given finite field. In section 6 we prove the following statement.
Theorem 1.5.
Let $G$ be a finite group, and let $k$ be a field of characteristic $p>2$. Assume that $p$ does not divide the order of $G$. There exists an abelian surface with a faithful action of $G$ over $k$ such that $A/G$ is birational to a $K3$ surface if and only if $G$ belongs to the Katsura list from Theorem 1.4, and one of the following conditions hold:
  1. (1)
    $G$ is $Q_{8},Q_{12},\operatorname{SL}_{2}({\mathbb{F}}_{3})$, or a cyclic group of order $2,3,4$, or $6$;
  2. (2)
    if $G$ contains a cyclic subgroup of order $n\in\{5,8,12\}$, then $A$ is supersingular, ${\mathbb{F}}_{p^{2}}\subset k$, and $p\not\equiv\pm 1\bmod n$.
To prove the theorem, we use that there is a rigid action of a finite group $G$ on an abelian variety $A$ if and only if there is a homomorphism from a certain semisimple algebra ${\mathbb{Q}}[G]^{\mathrm{rig}}$ to the endomorphism algebra $\operatorname{End}(A)\otimes{\mathbb{Q}}$ of $A$ (see Section 3). The algebra ${\mathbb{Q}}[G]^{\mathrm{rig}}$ is the rigid quotient of the group algebra ${\mathbb{Q}}[G]$. In Section 4 we obtain a classification of groups $G$ with a rigid action on an abelian surface. This result can be extracted from the results of Hwang [Hw], but we feel that in this particular case it is easier to give an independent proof.
Finally, in section 7 we study traces of Frobenius actions on Neron–Severi groups of supersingular generalized Kummer surfaces. A supersingular $K3$ surface $X_{21}$ over ${\mathbb{F}}_{p}$ with the zeta function $(1^{21},2)$ was constructed by Schuett [Sch12]. We give an independent construction on the assumption that $p\equiv 3\bmod 4$. We summarize Theorems 7.4 and 7.11 as follows.
Theorem 1.6.
There exist supersingular $K3$ surfaces over ${\mathbb{F}}_{q}$, where $q$ is an even power of $p$, with the following traces: $22,18,14,10,8,6,4,2,0$
There exist supersingular $K3$ surfaces over ${\mathbb{F}}_{q}$, where $q$ is an odd power of $p$, with the following traces: $20,18,14,10,8,6,2,0$
Note that in all our examples the traces are non-negative and even. It is natural to ask whether the trace is always non-negative and even for a supersingular $K3$ surface over ${\mathbb{F}}_{q}$.
The author thanks Yuri Zarhin and Matthias Schütt for interesting discussions and WonTae Hwang for his remarks on the paper. The author is a Simons-IUM contest winner and thanks its sponsors and jury.
2. Preliminaries
2.1. Notation
[…]
2.2. Central simple algebras over number fields
Recall some well-known facts on central simple algebras over number fields [MO, Section 28]. Let $L/K$ be an extension of number fields, and let ${\mathbb{H}}$ be a simple algebra with center $K$ such that $L\otimes{\mathbb{H}}$ is a matrix algebra over $L$, i.e., ${\mathbb{H}}$ represents an element of the Brauer group $\operatorname{Br}(L/K)$. Let $v$ be a prime of $K$. We denote the local invariant of ${\mathbb{H}}$ in $v$ by $\operatorname{inv}_{v}{\mathbb{H}}\in{\mathbb{Q}}/{\mathbb{Z}}$. A simple algebra over a number field $K$ is uniquely determined by its local invariants in the places of $K$ [MO, Remark 32.12].
Theorem 2.1.
Let ${\mathbb{H}}$ be a central simple algebra over $K$, and let $v$ be a prime of $K$. Let $L$ be an extension of $K$. Choose a prime $w$ of $L$ over $v$. Then ${\mathbb{H}}_{L}$ is a central simple algebra over $L$ with local invariant $[L_{w}:K_{v}]\operatorname{inv}_{v}{\mathbb{H}}$ in $w$.
Proof.
It follows from [MO, Theorem 31.9]. ∎
Let $K$ be a totally real extension of ${\mathbb{Q}}$ of even degree. We denote by ${\mathbb{H}}_{\infty}(K)$ the quaternion algebra with center $K$, and non-trivial invariants exactly in the real places of $K$. We denote by ${\mathbb{H}}_{p}$ the quaternion algebra with center ${\mathbb{Q}}$, and non-trivial invariants in the real place and in $p$.
Theorem 2.2.
Let ${\mathbb{H}}$ be a simple algebra of dimension $d^{2}$ over its center $K$. Let $L$ be an extension of $K$ of degree $d$. Then there is a homomorphism of $L$ to ${\mathbb{H}}$ if and only if $L\otimes_{K}{\mathbb{H}}$ is a matrix algebra.
Proof.
It follows from [MO, Theorem 28.5]. ∎
Example 2.3.
Let $L$ be a CM number field, and denote by $L^{\mathrm{real}}$ the totally real subfield of $L$. According to the previous theorem, there exists an injection $L\to{\mathbb{H}}_{\infty}(L^{\mathrm{real}})$.
Corollary 2.4.
  1. (1)
    Let $K$ be a real quadratic extension of ${\mathbb{Q}}$. There is a homomorphism of ${\mathbb{H}}_{p}$ to ${\mathbb{H}}_{\infty}(K)$ if and only if $p$ is not split in $K$.
  2. (2)
    Let $L$ be an imaginary quadratic extension of ${\mathbb{Q}}$. There is a homomorphism of ${\mathbb{H}}_{p}$ to $M(2,L)$ if and only if $p$ is not split in $L$.
Proof.
  1. (1)
    If $p$ is not split in $K$, then, according to Theorems 2.1 and 2.2, we have a homomorphism
    ${\mathbb{H}}_{p}\to{\mathbb{H}}_{p}\otimes K\cong{\mathbb{H}}_{\infty}(K).$
    Assume that $p$ is split in $K$, and that there is a homomorphism ${\mathbb{H}}_{p}\to{\mathbb{H}}_{\infty}(K)$. Then we get a homomorphism of the simple algebra ${\mathbb{H}}_{p}\otimes{\mathbb{Q}}_{p}$ to the sum of two matrix algebras
    ${\mathbb{H}}_{\infty}(K)\otimes{\mathbb{Q}}_{p}\cong M(2,{\mathbb{Q}}_{p})\oplus M(2,{\mathbb{Q}}_{p}).$
    A contradiction.
  2. (2)
    There is a homomorphism of ${\mathbb{H}}_{p}$ to $M(2,L)$ if and only if there is a homomorphism in the opposite direction of the centralizers of these algebras in $M(4,{\mathbb{Q}})$, i.e., $L\to{\mathbb{H}}_{p}$. According to Theorems 2.1 and 2.2, the last homomorphism exists if and only if $p$ is not split in $L$.
2.3. Abelian varieties over finite fields
Let $A$ be an abelian variety over $k$. The endomorphism ring of an abelian variety $\operatorname{End}(A)$ is finitely generated and torsion-free as ${\mathbb{Z}}$-module. Let
$\operatorname{End}^{\circ}(A)=\operatorname{End}(A)\otimes_{{\mathbb{Z}}}{\mathbb{Q}}.$
An element $\varphi\in\operatorname{End}(A)$ is called an isogeny if $\varphi$ is finite and surjective. For example, multiplication by $m\in{\mathbb{Z}}$ is an isogeny $[m]:A\to A$. Denote the kernel of $[m]$ by $A[m]$. This is a finite group scheme over $k$ of order $m^{2\dim A}$ [Mil08, Remark 7.3].
An abelian variety $A$ is simple if it does not contain nontrivial abelian subvarieties. Any abelian variety $A$ over $k$ is isogenous to a product of simple abelian varieties:
$A\to\prod_{i}A_{i}^{r_{i}},$
where $A_{i}$ are simple. This decomposition corresponds to a decomposition of $\operatorname{End}^{\circ}(A)$ into a product of simple algebras $\operatorname{Mat}(r_{i},\operatorname{End}^{\circ}(A_{i}))$. In particular, $\operatorname{End}^{\circ}(A)$ is a semisimple ${\mathbb{Q}}$-algebra [Mum70, IV.19, Corollaries 1 and 2].
Fix a prime number $\ell\neq p$. Let
$T_{\ell}(A)=\mathop{{\lim\limits_{\longleftarrow}}}\nolimits A[\ell^{r}]({\bar{k}}),\text{ and }V_{\ell}(A)=T_{\ell}(A)\otimes_{{\mathbb{Z}}_{\ell}}{\mathbb{Q}}_{\ell}$
be the $\ell$-th Tate module of $A$, and the corresponding vector space over ${\mathbb{Q}}_{\ell}$. It is known that $T_{\ell}(A)$ is a free ${\mathbb{Z}}_{\ell}$-module of rank $2\dim A$.
Let $k$ be a finite field ${\mathbb{F}}_{q}$. The Frobenius endomorphism $F_{A}:A\to A$ over $k$ is the morphism that is trivial on the topological space and raises functions to their $q$-th powers. The endomorphism $F_{A}$ acts on the Tate module by a semisimple linear transformation, which we denote by $F:T_{\ell}(A)\to T_{\ell}(A)$. The characteristic polynomial
$f_{A}(t)=\det(t-F)$
is called the Weil polynomial of $A$. It is a monic polynomial of degree $2\dim A$ with rational integer coefficients independent of the choice of $\ell$ (see [Mum70, Page 180, Theorem 4], and [Mum70, Page 205]).
The Frobenius endomorphism generates the center $K_{A}\subset\operatorname{End}^{\circ}(A)$. Tate proved that the isogeny class of an abelian variety over $k$ is determined by its characteristic polynomial, that is $f_{A}(t)=f_{B}(t)$ if and only if $A$ is isogenous to $B$ [Ta66]. Moreover, for any $\ell\neq p$ we have
$\operatorname{End}(A)\otimes_{{\mathbb{Z}}}{\mathbb{Z}}_{\ell}\cong\operatorname{End}_{F}(T_{\ell}A),$
where $\operatorname{End}_{F}(T_{\ell}A)$ is the ring of endomorphisms of $T_{\ell}(A)$ that commute with $F$. For any embedding $\sigma:K_{A}\to{\mathbb{C}}$ we have $|\sigma(F_{A})|=\sqrt{q}$ [Mum70, Page 206, Theorem 4]. The following lemma can be proved in the same way as [Mil08, IV.2.3].
Lemma 2.5.
If $f:A\to B$ is an isogeny, then $T_{\ell}(f):T_{\ell}(A)\to T_{\ell}(B)$ is injective, and the preimage $T=T_{\ell}(f)^{-1}(T_{\ell}(B))$ is an $F$-invariant submodule of $V_{\ell}(A)$. Conversely, if $T\subset V_{\ell}(A)$ is $F$-invariant ${\mathbb{Z}}_{\ell}$-submodule of finite rank such that $T_{\ell}(A)\subset T$, then there exists an abelian variety $B$ defined over $k$ and an $\ell$-isogeny $f:A\to B$ such that $T_{\ell}(f)$ induces an isomorphism $T\cong T_{\ell}(B)$.∎
Assume that $A$ is simple. In this case $K_{A}$ is a field. Let $L_{A}$ be the Galois envelope of $K_{A}$ over ${\mathbb{Q}}$, and let $P$ be the set of primes of $L_{A}$ over $p$. The slopes of $A$ are defined as the set
$\{\frac{v_{\mathfrak{p}}(F_{A})}{v_{\mathfrak{p}}(q)}|\mathfrak{p}\in P\},$
where $v_{\mathfrak{p}}$ is the normalized valuation. We say that $A$ is supersingular if all slopes of $A$ are equal to $1/2$.
Lemma 2.6.
Let $L$ be a Galois extension of ${\mathbb{Q}}$. Assume that $\sqrt{q}\not\in{\mathbb{Z}}$, and $p$ is unramified in $L$. Then there is no supersingular abelian variety $A$ such that $L_{A}=L$.
Proof.
Let $A$ be an abelian variety such that $L_{A}=L$. Then the slopes of $A$ are the fractions ${v_{\mathfrak{p}}(F_{A})}/{v_{\mathfrak{p}}(q)}$ with odd denominator. ∎
The abelian variety is ordinary, if the set of slopes of $A$ is $\{0,1\}$. It is known that $A$ is ordinary if and only if the order of the group $A[p]({\bar{k}})$ is equal to $p^{\dim A}$. A simple abelian surface $A$ is called mixed, if the set of slopes of $A$ is $\{0,1/2,1\}$.
In dimensions $1$ and $2$ Weil polynomials can be explicitly classified. We present this classification in part. For more details, see [MN02] and [Wa69].
Theorem 2.7.
[Wa69] Let $E$ be an elliptic curve over ${\mathbb{F}}_{q}$. Then $f_{E}(t)=t^{2}-bt+q$, where $|b|\leq 2\sqrt{q}$.
  1. (1)
    $E$ is ordinary if and only if $v_{p}(b)=0$. In this case
    $\operatorname{End}^{\circ}(E)\cong{\mathbb{Q}}[t]/f_{E}(t){\mathbb{Q}}[t].$
  2. (2)
    If $E$ is supersingular and $f_{E}$ is separable, then
    $\operatorname{End}^{\circ}(E)\cong{\mathbb{Q}}[t]/f_{E}(t){\mathbb{Q}}[t],$
    and $q$ and $b$ satisfy the following conditions:
    1. (a)
      $\sqrt{q}\not\in{\mathbb{Z}}$, and $b=0$;
    2. (b)
      $\sqrt{q}\in{\mathbb{Z}}$, $b=0$, and $p\not\equiv 1\bmod 4$;
    3. (c)
      $\sqrt{q}\in{\mathbb{Z}}$, $b=\pm\sqrt{q}$, and $p\not\equiv 1\bmod 3$;
    4. (d)
      $p=2$ or $p=3$, $\sqrt{q}\not\in{\mathbb{Z}}$, and $b=\pm\sqrt{pq}$.
  3. (3)
    If $E$ is supersingular and $f_{E}$ is not separable, then $\sqrt{q}\in{\mathbb{Z}}$ is an integer, $f_{E}(t)=(t\pm\sqrt{q})^{2}$, and $\operatorname{End}^{\circ}(E)\cong{\mathbb{H}}_{p}$.
In all these cases, there exists an elliptic curve over ${\mathbb{F}}_{q}$ with a given Weil polynomial.
Theorem 2.8.
[Wa69, Ru90, MN02] Let $A$ be a simple abelian surface over ${\mathbb{F}}_{q}$, where $q=p^{n}$. Then $f_{A}(t)=P_{A}(t)^{e}$, where $P_{A}\in{\mathbb{Z}}[t]$ is irreducible.
  1. (1)
    If $A$ is ordinary or mixed, then $e=1$, and
    $\operatorname{End}^{\circ}(A)\cong{\mathbb{Q}}[t]/f_{A}(t){\mathbb{Q}}[t].$
  2. (2)
    If $A$ is supersingular, then we have the following possibilities:
    1. (a)
      $e=1$, and $\operatorname{End}^{\circ}(A)\cong{\mathbb{Q}}[t]/f_{A}(t){\mathbb{Q}}[t]$. In this case $f_{A}(t)=t^{4}+a_{1}t^{3}+a_{2}t^{2}+a_{1}qt+q^{2}$ is irreducible, and the pair $(a_{1},a_{2})$ belongs to the following list:
      1. (i)
        $(0,0)$, $n$ is odd, and $p\neq 2$;
      2. (ii)
        $(0,0)$, $n$ is even, and $p\not\equiv 1\bmod 8$;
      3. (iii)
        $(0,q)$, and $n$ is odd;
      4. (iv)
        $(0,-q)$, $n$ is odd, and $p\neq 3$;
      5. (v)
        $(0,-q)$, $n$ is even, and $p\not\equiv 1\bmod 12$;
      6. (vi)
        (±$(\pm\sqrt{q},q)$, $n$ is even, and $p\not\equiv 1\bmod 5$;
      7. (vii)
        (±$(\pm\sqrt{5q},3q)$, $n$ is odd, and $p=5$;
      8. (viii)
        (±$(\pm\sqrt{2q},q)$, $n$ is odd, and $p=2$.
    2. (b)
      $e=2$, $n$ is odd, $P_{A}(t)=(t^{2}-q)^{2}$, and $\operatorname{End}^{\circ}(A)\cong{\mathbb{H}}_{\infty}({\mathbb{Q}}(\sqrt{p}))$;
    3. (c)
      $e=2$, $n$ is even, $P_{A}(t)=t^{2}-bt+q$, and one of the following conditions holds
      1. (i)
        $b=0$, and $p\equiv 1\bmod 4$;
      2. (ii)
        $b=\pm\sqrt{q}$, and $p\equiv 1\bmod 3$.
      In both cases $\operatorname{End}^{\circ}(A)$ is a quaternion algebra over ${\mathbb{Q}}[t]/P_{A}(t){\mathbb{Q}}[t]$.
In all these cases, there exists a simple abelian surface over ${\mathbb{F}}_{q}$ with a given Weil polynomial.
2.4. Diedonné modules of abelian varieties
Let $W(k)$ be the ring of Witt vectors over $k$. We denote by $\sigma:W(k)\to W(k)$ the lift of the Frobenius automorphism of $k$. The Diedonné ring $D_{k}$ is the ring generated over $W(k)$ by formal variables $F$ and $V$ such that $FV=VF=p$, and for any $a\in W(k)$ we have:
$Fa=\sigma(a)F,\text{ and }aV=V\sigma(a).$
There is a contravariant equivalence of categories $M(-)$ between finite group schemes over $k$ of $p$-power order and left $D_{k}$-modules of finite length [Pink, Theorem 28.3].
The $p^{r}$-torsion $A[p^{r}]$ of an abelian variety $A$ over $k$ is a finite group scheme, and the limit
[…]
is called the Diedonné module of $A$. The module $M(A)$ is free over $W(k)$ of rank $2\dim A$, and the cotangent space $T^{*}_{0}(A)$ of $A$ at zero is canonically isomorphic to $M(A)/FM(A)$ [Pink, Proposition 28.4]. We have the following analog of Lemma 2.5.
Lemma 2.9.
If $f:B\to A$ is an isogeny then, $M(f)$ is injective, and $M=M(f)^{-1}(M(B))$ is a Diedonné submodule of $M(A)\otimes{\mathbb{Q}}$.
Conversely, if $M\subset M(A)\otimes{\mathbb{Q}}$ is a Diedonné submodule of finite ${\mathbb{Z}}_{p}$-rank such that $M(A)\subset M$, then there exists an abelian variety $B$ over $k$ and a $p$-isogeny $f:B\to A$ such that $M(f)$ induces an isomorphism $M(B)\cong M$.∎
Let $r,s\in{\mathbb{N}}$ be natural numbers. We say that a Diedonné module $M$ is pure of slope $r/s$, if there exists a submodule $M^{\prime}\subset M$ such that $M^{\prime}\otimes{\mathbb{Q}}\cong M\otimes{\mathbb{Q}}$, and $F^{s}(M^{\prime})=p^{r}M^{\prime}$.
Theorem 2.10.
[Ma63, §4][DO12, Theorem 1.3] Let $M$ be a Diedonné module. Then
$M\otimes{\mathbb{Q}}\cong\oplus_{\lambda}M_{\lambda}\otimes{\mathbb{Q}},$
where each $M_{\lambda}\subset M$ is pure of slope $\lambda\in{\mathbb{Q}}$.
We say that the set $\{\lambda\in{\mathbb{Q}}|M_{\lambda}\neq 0\}$ is the set of slopes of $M$. If $k$ is a finite field, and $M=M(A)$ is the Diedonné module of an abelian variety $A$ over $k$, then by [Ma63, Theorem 4.1] the slopes of $M$ and slopes of $A$ coincide. This result motivates a more general definition. An abelian variety $A$ over a perfect field $k$ is supersingular, if all slopes of $M(A)$ are equal to $1/2$, and $A$ is ordinary, if the set of slopes of $M(A)$ is $\{0,1\}$. The following result is due to Tate, Shioda, Deligne, and Oort.
Theorem 2.11.
[Oort, Theorem 4.2] Let $k$ be a perfect field of characteristic $p$, and let $E$ be a supersingular elliptic curve over a finite field ${\mathbb{F}}_{q}\subset k$. Then any supersingular abelian surface $A$ over $k$ is isogenous to $E^{2}$ over ${\bar{k}}$. In particular, $\operatorname{End}^{\circ}(A)$ is a subalgebra of $M(2,{\mathbb{H}}_{p})$.
The structure of the Diedonné module of an ordinary abelian variety is well known.
Proposition 2.12.
Let $A$ be an ordinary abelian variety over $k$. Then $M(A)\cong M\oplus M^{*}$, where $M$ is a Diedonné module of slope $1$ and rank $\dim(A)$ over $W(k)$, and $T^{*}_{0}(A)\cong M/pM.$
Proof.
According to [Pink, Proposition 15.4], the group scheme $A[p^{n}]$ is a direct sum of its étale part $X^{*}_{n}$, and the local part $X_{n}$. Since $A$ is ordinary, the order of $X^{*}_{n}$ is $p^{n\dim A}$; therefore, the order of $X_{n}$ is also $p^{n\dim A}$. It follows that
[…]
are free $W(k)$-modules of rank $\dim A$, and $M(A)\cong M\oplus M^{*}$. The Frobenius endomorphism is invertible on the étale part $M^{*}$; therefore, according to [Pink, Proposition 28.4],
$T^{*}_{0}(A)\cong M(A)/FM(A)\cong M/pM.$
The proposition is proved. ∎
2.5. Finite group actions on abelian varieties
Let $G$ be a finite group, and an action of $G$ on an abelian variety $A$ is given by a homomorphism $G\to\operatorname{End}(A)^{*}$. The induced action on $T^{*}_{0}(A)$ gives a homomorphism
$G\to\operatorname{GL}_{d}(k),$
where $d$ is the dimension of $A$.
Lemma 2.13.
Let $g\in G$ be an element of order $r$. If the image of $g$ in $\operatorname{GL}_{d}(k)$ is trivial, then $r$ is a power of $p$.
Proof.
We may assume that $r\neq p$ is the prime and that the action of $g$ on $T^{*}_{0}(A)$ is trivial. The group scheme corresponding to the Diedonné module $M(A)/FM(A)$ is a subscheme of $\ker(g-1)$; therefore, the eigenvalues of $M(g)-1$ are not $p$-adic units. Since the norm
$N_{{\mathbb{Q}}(\zeta_{r})/{\mathbb{Q}}}(\zeta_{r}-1)=r,$
is a $p$-adic unit, we get a contradiction with the fact that the eigenvalues of $M(g)$ are roots of unity of order $r\neq p$. ∎
Define a decreasing filtration on $G$ as follows: $G=G_{0}$, and for $s>0$
$G_{s}=\ker(G\to\operatorname{Aut}(M(A)/F^{s}M(A))).$
By definition, $G_{0}/G_{1}$ is a subgroup of $\operatorname{GL}_{2}(k)$.
Lemma 2.14.
Let $s>0$. We have an inclusion $G_{s}/G_{s+1}\to\operatorname{Hom}(T^{*}_{0}A,T^{*}_{0}A)$. In particular, the quotient $G_{s}/G_{s+1}$ is abelian.
Proof.
If $g$ is an element of $G_{s}$, then for any $v\in M(A)/F^{s+1}M(A)$ we have $g(v)=v+F^{s}v_{g}$ for some unique $v_{g}\in M(A)/FM(A)$. Moreover, if $v\in FM(A)$, then $v_{g}=0$. Since $F$ is injective on $M(A)$, the morphism $F^{s}$ induces an isomorphism
$M(A)/FM(A)\to F^{s}M(A)/F^{s+1}M(A).$
Define the morphism
$\alpha_{s}:G_{s}\to\operatorname{Hom}(M(A)/FM(A),M(A)/FM(A))$
as follows: lift $\bar{v}\in M(A)/FM(A)$ to some $v\in M(A)/F^{s+1}M(A)$, and put $\alpha_{s}(g)(\bar{v})=v_{g}$. Clearly, $\alpha_{s}(g)=0$ if and only if $g(v)=v$ for all $v\in M(A)/F^{s+1}M(A)$, that is, $g\in G_{s+1}$. ∎
Lemma 2.15.
Let $Q$ be a $p$-Sylow subgroup of $G$. Then $G_{1}$ is normal in $Q$, and $Q/G_{1}$ is annihilated by $p$.
Proof.
The group $G_{1}$ is normal in $G$; therefore, it is normal in $Q$ as well. The quotient $Q/G_{1}$ is isomorphic to a subgroup of $\operatorname{GL}_{2}({\mathbb{F}}_{q})$, and the $p$-Sylow subgroup of $\operatorname{GL}_{2}({\mathbb{F}}_{q})$ is annihilated by $p$. ∎
Corollary 2.16.
If the $p$-th Sylow subgroup $Q$ of $G$ is not annihilated by $p$, then $G_{1}$ is non-trivial.
Example 2.17.
Let $p=2$, and let $G=\operatorname{SL}_{2}({\mathbb{F}}_{5})$. The $2$-nd Sylow subgroup of $G$ is $Q_{8}$; therefore, the filtration given by subgroups $G_{i}$ is non-trivial.
3. Rigid actions on abelian varieties
3.1. An equivalence of categories.
Let $\mathcal{C}$ be an isogeny class of abelian varieties over $k$, and let $G$ be a finite group. Denote by $\mathcal{C}_{G}$ the category of abelian varieties from $\mathcal{C}$ with an action of $G$ that fixes the group law. Recall that such an action on an abelian variety $A$ is given by a homomorphism $G\to\operatorname{End}(A)^{*}$. We define the $\operatorname{Hom}$ group as follows:
$\operatorname{Hom}_{\mathcal{C}_{G}}(A,B)=\operatorname{Hom}^{\circ}(A,B)^{G}.$
Let $\mathcal{D}_{G}$ be the following category. Objects of $\mathcal{D}_{G}$ are group homomorphisms $G\to\operatorname{End}^{\circ}(A)^{*}$, where $A$ in $\mathcal{C}$, and a morphism from $i_{A}:G\to\operatorname{End}^{\circ}(A)^{*}$ to $i_{B}:G\to\operatorname{End}^{\circ}(B)^{*}$ is a $\psi\in\operatorname{Hom}^{\circ}(A,B)$ that fits the diagram:
[…]
There is a natural functor from $\mathcal{C}_{G}$ to $\mathcal{D}_{G}$ sending an abelian variety $A$ with a $G$-action to the induced homomorphism $G\to\operatorname{End}^{\circ}(A)^{*}$.
Proposition 3.1.
The natural functor from $\mathcal{C}_{G}$ to $\mathcal{D}_{G}$ is an equivalence of categories.
Proof.
Clearly, the functor is full and faithful. We have to prove that the functor is essentially surjective. Let $A^{\prime}$ be an abelian variety in $\mathcal{C}$, and let $i_{A^{\prime}}:G\to\operatorname{End}^{\circ}(A^{\prime})^{*}$ be a homomorphism. For any $\ell\neq p$ we define a $G$-invariant submodule $T_{\ell}$ by the formula
$T_{\ell}=\sum_{g\in G}gT_{\ell}(A^{\prime})\subset V_{\ell}(A^{\prime}).$
We claim that $T_{\ell}=T_{\ell}(A^{\prime})$ for almost all $\ell\neq p$. Indeed, the order $\operatorname{End}(A^{\prime})\otimes{\mathbb{Z}}_{\ell}$ is maximal at almost all $\ell$, and is equal to the order of integral elements in $\operatorname{End}^{\circ}(A^{\prime})\otimes{\mathbb{Q}}_{\ell}$ [MO, Theorem 12.8]. Since the image of $G$ in $\operatorname{End}^{\circ}(A^{\prime})$ consists of integral elements, for such $\ell$ we have
$GT_{\ell}(A^{\prime})\subset T_{\ell}(A^{\prime}).$
We have a finite number of primes $\ell\neq p$ such that $T_{\ell}\neq T_{\ell}(A^{\prime})$. Now we apply Lemma 2.5 to all such $T_{\ell}$ and get an isogeny $A^{\prime}\to A^{\prime\prime}$ such that $T_{\ell}\cong T_{\ell}(A^{\prime\prime})$ for all $\ell\neq p$. Finally, we put
$M_{p}=\sum_{g\in G}gM(A^{\prime})\subset M(A^{\prime})\otimes{\mathbb{Q}}_{p}.$
According to Lemma 2.9, there exists an isogeny $A\to A^{\prime\prime}$ such that $T_{\ell}\cong T_{\ell}(A)$, and $M_{p}\cong M(A)$. In particular,
$i_{A}(G)\subset\operatorname{End}(A)\otimes{\mathbb{Z}}_{\ell}$
for all $\ell$. Consider the natural inclusion $\iota:\operatorname{End}(A)\to T_{G}$ to the lattice generated by $\operatorname{End}(A)$ and the image of $i_{A}(G)$. For all primes $\ell$ the localization $\iota\otimes{\mathbb{Z}}_{\ell}$ is an isomorphism; this implies that $T_{G}=\operatorname{End}(A)$, and $G\subset\operatorname{End}(A)$. We proved that there exists an abelian variety $A$ in $\mathcal{C}_{G}$, and the diagram $A^{\prime}\to A^{\prime\prime}\leftarrow A$ induces an isomorphism in $\mathcal{D}_{G}$. ∎
Corollary 3.2.
If there exists an abelian variety $A$ with a $G$-action over $k$, then there exists an abelian variety $B$ of the same dimension over a finite field endowed with a $G$-action and an injective homomorphism $\operatorname{End}^{\circ}(A)\to\operatorname{End}^{\circ}(B)$ such that the $G$-action is given by the composition
$G\to\operatorname{End}^{\circ}(A)^{*}\to\operatorname{End}^{\circ}(B)^{*}.$
If $A$ is supersingular, then $B$ is supersingular as well.
Proof.
Let ${\mathbb{F}}_{q}$ be the algebraic closure of ${\mathbb{F}}_{p}$ in $k$. There exists a smooth connected affine variety $U$ over ${\mathbb{F}}_{q}$ such that $k$ is a field of functions on $U$. If we make $U$ smaller, then there exists a smooth abelian scheme $\mathcal{A}$ over $U$ such that $A$ is the general fiber of $\mathcal{A}$. Let $B$ be a special fiber of $\mathcal{A}$ over a finite extension of ${\mathbb{F}}_{q}$. Then there exists a natural injection [CMLP, Section 1.8.4.1]:
$\operatorname{End}(A)\to\operatorname{End}(B).$
Therefore, $G$ acts on $B$.
The Newton polygon of $A$ lies not higher than the Newton polygon of $B$ by the Grothendieck–Katz specialization theorem [Ka78]; therefore, if $A$ is supersingular, then $B$ is also supersingular. ∎
Proposition 3.3.
The action of $G$ on $A$ is rigid if and only if any non-trivial $g\in G$ has only finite number of fixed points.
Proof.
The eigenvalues of the action of $g\in G$ on $V_{\ell}(A)$ are primitive roots of unity if and only if $\ker(1-V_{\ell}(g)^{r})\subset V_{\ell}(A)$ is trivial for any divisor $r$ of the order of $g$. Let $B$ be the connected component of identity in $\ker(1-g^{r})$. Then $B$ is an abelian subvariety, and $g^{r}$ has only a finite number of fixed points if and only if $\dim B=0$, and if and only if $V_{\ell}(B)=0$. We have to prove that $V_{\ell}(B)=\ker(V_{\ell}(1-g^{r}))$. Clearly, $V_{\ell}(B)\subset\ker(V_{\ell}(1-g^{r}))$. We will prove that any $v\in\ker(V_{\ell}(1-g^{r}))$ belongs to $V_{\ell}(B)$. Since $V_{\ell}(A)=T_{\ell}(A)\otimes{\mathbb{Q}}_{\ell}$, we have $v=v^{\prime}\otimes\frac{1}{\ell^{a}}$ for some natural $a$. The vector $v^{\prime}$ is represented by a sequence $v_{j}\in A[\ell^{j}]$ such that $v_{j}=\ell v_{j+1}$, and $g^{r}(v_{j})=v_{j}$. Since the quotient $\ker(1-g^{r})/B$ is finite, $v_{j}\in B$ for large $j$; therefore, $v\in V_{\ell}(B)$. ∎
Corollary 3.4.
If $A$ is simple over $k$, then any faithful action of a finite group on $A$ is rigid.
Denote by $\mathcal{C}_{G}^{\mathrm{rig}}$ the full subcategory in $\mathcal{C}_{G}$ of abelian varieties with a rigid action of $G$. The homomorphism $i_{A}:G\to\operatorname{End}^{\circ}(A)^{*}$ induces a representation on $V_{\ell}(A)$. We say that $i_{A}$ is rigid if this representation is without fixed points. Let $\mathcal{D}_{G}^{\mathrm{rig}}$ be the full subcategory of rigid objects in $\mathcal{D}_{G}$. According to Proposition 3.1, $\mathcal{C}_{G}^{\mathrm{rig}}$ and $\mathcal{D}_{G}^{\mathrm{rig}}$ are equivalent.
Let ${\mathbb{H}}$ be a semisimple algebra over ${\mathbb{Q}}$. Denote by $\mathcal{C}_{\mathbb{H}}$ the following category. Objects of $\mathcal{C}_{\mathbb{H}}$ are homomorphisms of ${\mathbb{Q}}$-algebras ${\mathbb{H}}\to\operatorname{End}^{\circ}(A)$, where $A$ is in $\mathcal{C}$, and a morphism from $h_{A}:{\mathbb{H}}\to\operatorname{End}^{\circ}(A)$ to $h_{B}:{\mathbb{H}}\to\operatorname{End}^{\circ}(B)$ is an element $\psi\in\operatorname{Hom}^{\circ}(A,B)$ that fits in the diagram:
[…]
The rational group algebra ${\mathbb{Q}}[G]$ of $G$ is isomorphic to the sum of simple algebras corresponding to irreducible representations $V$ over ${\mathbb{Q}}$:
${\mathbb{Q}}[G]=\oplus_{V}{\mathbb{H}}_{V}.$
Define the rigid group algebra as the sum over representations without fixed points (WFP):
${\mathbb{Q}}[G]^{\mathrm{rig}}:=\oplus_{V\text{ is WFP}}{\mathbb{H}}_{V}.$
A homomorphism $i_{A}$ from $G$ to $\operatorname{End}^{\circ}(A)^{*}$ induces the natural homomorphism of ${\mathbb{Q}}$-algebras $h_{A}:{\mathbb{Q}}[G]\to\operatorname{End}^{\circ}(A)$. If $i_{A}$ is rigid, then $h_{A}$ is the composition of the natural projection ${\mathbb{Q}}[G]\to{\mathbb{Q}}[G]^{\mathrm{rig}}$ and a homomorphism
$h_{A}^{\mathrm{rig}}:{\mathbb{Q}}[G]^{\mathrm{rig}}\to\operatorname{End}^{\circ}(A).$
For ${\mathbb{H}}={\mathbb{Q}}[G]^{\mathrm{rig}}$ this gives an equivalence of $\mathcal{D}_{G}^{\mathrm{rig}}$ and $\mathcal{C}_{\mathbb{H}}$. We proved the main result of this section.
Theorem 3.5.
Let $G$ be a group with a representation without fixed points, and let ${\mathbb{H}}={\mathbb{Q}}[G]^{\mathrm{rig}}$. Then the categories $\mathcal{C}_{G}^{\mathrm{rig}}$ and $\mathcal{C}_{\mathbb{H}}$ are equivalent.∎
Lemma 3.6.
Let $G=C_{n}$ be a cyclic group of order $n$.
  1. (1)
    We have ${\mathbb{Q}}[G]^{\mathrm{rig}}\cong{\mathbb{Q}}(\zeta_{n})$.
  2. (2)
    Let $g$ be a generator of $G$. If $A$ is an abelian variety with a rigid action of $G$, then all primitive roots of unity of order $n$ are eigenvalues of the action of $g$ on $M(A)$ and on $V_{\ell}(A)$ for all $\ell\neq p$.
Proof.
Part $(1)$ follows from the isomorphism of algebras
${\mathbb{Q}}[G]\cong{\mathbb{Q}}\oplus{\mathbb{Q}}(\zeta_{n}).$
According to Theorem 3.5, if the action of $G$ on $A$ is rigid then, ${\mathbb{Q}}(\zeta_{n})$ is a subalgebra of $\operatorname{End}^{\circ}(A)$. By [Ri76, Theorem 2.1.1], $V_{\ell}(A)$ is a free module over ${\mathbb{Q}}(\zeta_{n})\otimes{\mathbb{Q}}_{\ell}$, where $\zeta_{n}$ acts $g$. Hence, $V_{\ell}(A)\otimes\bar{\mathbb{Q}}_{\ell}$ is a free module over
${\mathbb{Q}}(\zeta_{n})\otimes\bar{\mathbb{Q}}_{\ell}\cong\oplus_{\zeta}\bar{\mathbb{Q}}_{\ell},$
where the sum is over all immersions $\varphi_{\zeta}:{\mathbb{Q}}(\zeta_{n})\to\bar{\mathbb{Q}}_{\ell}$. Any such immersion is uniquely determined by the primitive root of unity of order $n$: $\zeta=\varphi_{\zeta}(\zeta_{n})$. It follows that,
$V_{\ell}(A)\otimes\bar{\mathbb{Q}}_{\ell}\cong\oplus_{\zeta}V_{\zeta},$
where $g$ acts on $V_{\zeta}$ with eigenvalue $\zeta$. The same argument can be applied to the Diedonné module $M(A)$. The lemma is proved. ∎
Corollary 3.7.
Let $G$ be a cyclic group of order $n$, and let ${\mathbb{H}}={\mathbb{Q}}(\zeta_{n})$ be the cyclotomic field. Then there is an equivalence of categories $\mathcal{C}_{G}$ and $\mathcal{C}_{\mathbb{H}}$.
Remark 3.8.
The equivalences of categories $\mathcal{C}_{G}$ and $\mathcal{C}_{\mathbb{H}}$ from the previous corollary are parametrized by identifications of $G$ and the group of $n$-th roots of unity.
Corollary 3.9.
There exists an abelian variety with a rigid action of $G$ in the isogeny class of an abelian variety $A$, if and only if there exists a homomorphism of ${\mathbb{Q}}$-algebras ${\mathbb{Q}}[G]^{\mathrm{rig}}\to\operatorname{End}^{\circ}(A).$
Corollary 3.10.
If there exists a supersingular abelian surface $A$ with a rigid action of $G$ over a perfect field $k$, then there exists a supersingular abelian surface with a rigid action of $G$ over ${\mathbb{F}}_{p^{2}}$. In particular, $G/G_{1}$ is a subgroup of $\operatorname{GL}_{2}({\mathbb{F}}_{p^{2}})$.
Proof.
Let $E$ be a supersingular elliptic curve over ${\mathbb{F}}_{p^{2}}$ such that $\operatorname{End}^{\circ}(E)\cong{\mathbb{H}}_{p}$. According to Theorem 2.11, $\operatorname{End}^{\circ}(A)$ is a subalgebra of $\operatorname{End}^{\circ}(E^{2})$. We get a homomorphism $G\to\operatorname{End}^{\circ}(E^{2})^{*}$. By Theorem 3.5, there exists an isogeny from $E^{2}$ to an abelian surface over ${\mathbb{F}}_{p^{2}}$ with a rigid action of $G$. ∎
3.2. Finite subgroups of endomorphism algebras.
The binary dihedral group $Q_{4n}$ of order $4n$ is a group generated by two elements $a$ and $b$ with the following relations:
$a^{n}=b^{2},\quad b^{4}=1,\quad aba=b.$
Lemma 3.11.
We have ${\mathbb{Q}}[Q_{8}]^{\mathrm{rig}}\cong{\mathbb{H}}_{2}$, ${\mathbb{Q}}[Q_{12}]^{\mathrm{rig}}\cong{\mathbb{H}}_{3}$, and for $n>3$
${\mathbb{Q}}[Q_{4n}]^{\mathrm{rig}}\cong{\mathbb{H}}_{\infty}({\mathbb{Q}}(\zeta_{2n})^{\mathrm{real}}).$
Proof.
Let $4n=2^{r}m$, where $m$ is odd. For any divisor $d$ of $m$ the element $a^{2n/d}$ generates a subgroup of order $d$, and the quotient of $Q_{4n}$ by this subgroup is isomorphic to $Q_{4n/d}$. This quotient corresponds to the following representation ${\mathbb{H}}(n,d)$. Let ${\mathbb{H}}_{\mathbb{R}}$ be the usual real quaternion algebra generated by $i$ and $j$ over its center ${\mathbb{R}}$. We generate a subring ${\mathbb{H}}(n,d)$ of ${\mathbb{H}}_{\mathbb{R}}$ by ${\mathbb{Q}}({\zeta_{2n/d}})\subset{\mathbb{C}}\cong{\mathbb{R}}[i]$, and $j$. In this representation, $a$ acts as a multiplication of $\zeta_{2n/d}$, and $b$ acts as $j$.
The quotient of $Q_{4n}$ by the subgroup generated by $b^{2}$ is $D_{2n}$. We find that the group algebra ${\mathbb{Q}}[Q_{2n}]$ has the following decomposition:
${\mathbb{Q}}[Q_{2n}]=\oplus_{d|m}{\mathbb{H}}(n,d)\oplus{\mathbb{Q}}[D_{2n}],$
due to the formula
$\dim{\mathbb{Q}}[Q_{2n}]=4n=2n+2\sum_{d|m}\varphi(2n/d)=\dim{\mathbb{Q}}[D_{2n}]+\sum_{d|m}\dim{\mathbb{H}}(n,d).$
The only summand corresponding to a faithful representation is ${\mathbb{H}}(n,1)$. This representation is rigid because $b^{2}$ acts as $-1$, and $a$ acts as $\zeta_{2n}$. Finally, ${\mathbb{H}}(2,1)\cong{\mathbb{H}}_{2}$, ${\mathbb{H}}(3,1)\cong{\mathbb{H}}_{3}$, and if $n>3$, then
${\mathbb{H}}(n,1)\cong{\mathbb{H}}_{\infty}({\mathbb{Q}}(\zeta_{2n})^{\mathrm{real}}).$
The lemma is proved. ∎
We need the following classic result of Burnside (see [Wo11, Section 5.3]).
Theorem 3.12.
[Bu] Let $G$ be a group with a representation without fixed points. Then all odd Sylow subgroups of $G$ are cyclic, and the even Sylow subgroup is either cyclic, or isomorphic to $Q_{2^{r}}$.
We say that a cyclic group of order $n$ is small, if $n\in\{2,3,4,5,6,8,10,12\}$.
Corollary 3.13.
Let $A$ be an abelian surface with a rigid action by a finite group $G$. Then the order of $G$ divides $240$, and any cyclic subgroup of $G$ is small.
Proof.
A cyclic subgroup of $G$ generates a commutative subalgebra of $\operatorname{End}^{\circ}(A)$ of dimension not greater than $4$; therefore, any cyclic subgroup is small. Moreover, since for an odd prime $p$ the $p$-Sylow subgroup $G^{(p)}$ is cyclic, we have $p\leq 5$, and $G^{(p)}\cong C_{p}$. We showed that the order of $G$ is $2^{r}m$, where $m$ divides $15$. If the $2$-Sylow subgroup is commutative, then, as before, $r\leq 3$. Otherwise, the subalgebra generated by the $2$-Sylow subgroup of $G$ contains ${\mathbb{Q}}[Q_{2^{r}}]^{\mathrm{rig}}$ as a direct summand. According to Lemma 3.11, $r\leq 4$. ∎
4. Rigid and symplectic actions on abelian surfaces
In this section we classify finite groups with a rigid and symplectic action on an abelian surface over $k$. Let $r$ be a power of an odd prime $p$. Denote by $\operatorname{ESL}_{2}({\mathbb{F}}_{r})$ the subgroup of $\operatorname{SL}_{2}({\mathbb{F}}_{r^{2}})$ generated by $\operatorname{SL}_{2}({\mathbb{F}}_{r})$ and an element given by the diagonal matrix
$\left(\begin{array}[]{c c}x&0\\ 0&x^{-1}\\ \end{array}\right),$
where $x\in{\mathbb{F}}_{r^{2}}\setminus{\mathbb{F}}_{r}$, and $x^{2}$ generates ${\mathbb{F}}_{r}^{*}$.
Remark 4.1.
The group $\operatorname{ESL}_{2}({\mathbb{F}}_{3})$ is the binary octahedral group.
Theorem 4.2 (Dickson).
[GT, Chapter 3, Theorem 6.17] Let $G$ be a finite subgroup of $\operatorname{SL}_{2}({\bar{k}})$. If the order of $G$ is relatively prime to $p$, then $G$ is isomorphic to one of the following groups:
  1. (1)
    a cyclic group $C_{n}$;
  2. (2)
    a binary dihedral group $Q_{2n}$;
  3. (3)
    $\operatorname{SL}_{2}({\mathbb{F}}_{3})$, $\operatorname{ESL}_{2}({\mathbb{F}}_{3})$, or $\operatorname{SL}_{2}({\mathbb{F}}_{5})$.
Suppose that $p$ divides the order of $G$, and $Q$ is a Sylow $p$–subgroup of $G$. Then one of the following cases occurs:
  1. (1)
    $G/Q$ is cyclic, and $Q$ is commutative and annihilated by $p$;
  2. (2)
    $G\cong\operatorname{SL}_{2}({\mathbb{F}}_{q})$, where $q$ is a power of $p$;
  3. (3)
    $p=2$, and $G$ is a dihedral group $D_{n}$ of order $2n$;
  4. (4)
    $p=3$, and $G\cong\operatorname{SL}_{2}({\mathbb{F}}_{5})$;
  5. (5)
    $\operatorname{ESL}_{2}({\mathbb{F}}_{q})$, where $q$ is a power of $p$.
We now state the main classification result of the section.
Theorem 4.3.
Let $G$ be a finite group with a rigid and symplectic action on an abelian surface over $k$. Then $G$ is one of the following groups.
  1. (1)
    $G$ is a small cyclic group of order $n$, and ${\mathbb{Q}}[G]^{\mathrm{rig}}\cong{\mathbb{Q}}(\zeta_{n})$;
  2. (2)
    $G$ is a binary dihedral group $Q_{4n}$ of order $4n$, where $2\leq n\leq 6$
    $G$ $Q_{8}$ $Q_{12}$ $Q_{16}$ $Q_{20}$ $Q_{24}$
    ${\mathbb{Q}}[G]^{\mathrm{rig}}$ ${\mathbb{H}}_{2}$ ${\mathbb{H}}_{3}$ ${\mathbb{H}}_{\infty}({\mathbb{Q}}(\sqrt{2}))$ ${\mathbb{H}}_{\infty}({\mathbb{Q}}(\sqrt{5}))$ ${\mathbb{H}}_{\infty}({\mathbb{Q}}(\sqrt{3}))$
  3. (3)
    $G$ is $\operatorname{SL}_{2}({\mathbb{F}}_{3})$, $\operatorname{ESL}_{2}({\mathbb{F}}_{3})$, or $\operatorname{SL}_{2}({\mathbb{F}}_{5})$
    $G$ $\operatorname{SL}_{2}({\mathbb{F}}_{3})$ $\operatorname{ESL}_{2}({\mathbb{F}}_{3})$ $\operatorname{SL}_{2}({\mathbb{F}}_{5})$
    ${\mathbb{Q}}[G]^{\mathrm{rig}}$ ${\mathbb{H}}_{2}$ ${\mathbb{H}}_{\infty}({\mathbb{Q}}(\sqrt{2}))$ ${\mathbb{H}}_{\infty}({\mathbb{Q}}(\sqrt{5}))$
  4. (4)
    $p=5$, and $G$ is $\operatorname{ESL}_{2}({\mathbb{F}}_{5})$, or $C_{5}\rtimes C_{8}$
    $G$ $\operatorname{ESL}_{2}({\mathbb{F}}_{5})$ $C_{5}\rtimes C_{8}$
    ${\mathbb{Q}}[G]^{\mathrm{rig}}$ $M(2,{\mathbb{H}}_{5})$ $M(2,{\mathbb{H}}_{5})$
  5. (5)
    $p=3$, and $G$ is $C_{3}\rtimes C_{8}$, $C_{3}\times Q_{8}$, or $C_{3}\rtimes Q_{16}$
    $G$ $C_{3}\rtimes C_{8}$ $C_{3}\times Q_{8}$ $C_{3}\rtimes Q_{16}$
    ${\mathbb{Q}}[G]^{\mathrm{rig}}$ $M(2,{\mathbb{Q}}[\zeta_{4}])$ $M(2,{\mathbb{Q}}[\zeta_{3}])$ $M(2,{\mathbb{H}}_{3})$
  6. (6)
    $p=2$, and $G$ is $C_{3}\rtimes C_{8}$, or $C_{3}\times Q_{8}$
    $G$ $C_{3}\rtimes C_{8}$ $C_{3}\times Q_{8}$
    ${\mathbb{Q}}[G]^{\mathrm{rig}}$ $M(2,{\mathbb{Q}}[\zeta_{4}])$ $M(2,{\mathbb{Q}}[\zeta_{3}])$
Remark 4.4.
In cases $(4),(5)$, and $(6)$ the algebra ${\mathbb{Q}}[G]^{\mathrm{rig}}$ is a subalgebra of $\operatorname{Mat}_{2}({\mathbb{H}}_{p})$. According to Theorem 2.2, and Corollary 3.9, there is a rigid action of the group $G$ on some supersingular abelian surface. In Section 6 we study rigid and symplectic actions in other cases.
Lemma 4.5.
If $p\geq 5$, then $G$ is a subgroup of $\operatorname{SL}_{2}(k)$.
Proof.
According to Lemma 2.13, any element of order prime to $p$ has a non-trivial image in $\operatorname{SL}_{2}(k)$. We have to prove the lemma for an element $g$ of order $5$ and $p=5$. The action of $g$ on the Diedonné module $M(A)$, induces an action of ${\mathbb{Z}}[\zeta_{5}]$ on $M(A)$. Note that ${\mathbb{Q}}_{5}[\zeta_{5}]$ is totally ramified over ${\mathbb{Q}}_{5}$; thus $M(A)$ is a free module of rank $1$ over the DVR $W(k)\otimes{\mathbb{Z}}_{5}[\zeta_{5}]$, and $1-g$ acts as a uniformiser. It follows that the action of $1-g$ on $M(A)/5M(A)$ is nilpotent with one-dimensional kernel. The cotangent space $T_{0}^{*}(A)$ is a two-dimensional quotient of $M(A)/5M(A)$; therefore, the action of $g$ on $T_{0}^{*}(A)$ is non-trivial. ∎
We now recall some basic facts on group extensions [GT]. Let $C$ be a normal subgroup of a group $E$. We say that $E$ is an extension of $C$ by $S=E/C$. If $C$ is abelian, then there is a natural action of $S$ on $C$, and equivalence classes of extensions with such an action correspond to elements of $H^{2}(S,C)$. In particular, if $H^{2}(S,C)$ is trivial, then $E\cong C\rtimes S$ is a semidirect product.
Lemma 4.6.
Assume that $G_{1}$ is abelian and that $\bar{G}=G/G_{1}$ has an abelian normal subgroup $H$ of prime to $p$ order. Then $H^{2}(\bar{G},G_{1})=H^{2}(\bar{G}/H,G_{1}^{H})$.
Proof.
Since $G_{1}$ is a $p$-group, $H^{1}(H,G_{1})=H^{2}(H,G_{1})=0$. The lemma follows from the inflation-restriction exact sequence:
$0\to H^{2}(\bar{G}/H,G_{1}^{H})\to H^{2}(\bar{G},G_{1})\to H^{2}(H,G_{1}).\qed$
Lemma 4.7.
The following groups do not have rigid representations: $D_{3}$, $D_{5}$, $C_{5}\rtimes C_{4}$, $D_{3}\times Q_{8}$, $C_{4}.A_{4}$, and $C_{4}.A_{5}$.
Proof.
The groups $D_{3}$, $D_{5}$, and $C_{5}\rtimes C_{4}$ have only one faithful representation, but the trace of an element of order $2$ is zero; therefore, the representations are not rigid. The groups $D_{3}\times Q_{8}$, $C_{4}.A_{4}$, and $C_{4}.A_{5}$ contain a subgroup $C_{2}^{2}$, and do not have rigid representations according to Remark 1.2.∎
Proof of the Theorem 4.3.
According to Lemma 3.13, the order of $G$ divides $240$, and any cyclic subgroup is small. If $p>5$, Lemma 4.5 and the Dickson Theorem 4.2 give first three cases. If $p=5$, then according to Lemma 4.5 and Dickson Theorem 4.2 we get that $G$ could be a group from cases $(1-3)$, $\operatorname{ESL}_{2}({\mathbb{F}}_{5})$, or an extension of $C_{5}$ by a cyclic group of order dividing $48$. It is straightforward to check that such extensions without non-small cyclic subgroups are either cyclic or belong to the following list: $D_{5}=C_{5}\rtimes C_{2}$, $Q_{20}$, $C_{5}\rtimes C_{4}$, and $C_{5}\rtimes C_{8}$. The groups $D_{5}$, and $C_{5}\rtimes C_{4}$ have no rigid representations according to Lemma 4.7.
Let $p=3$. Let $G_{1}$ be a normal $p$-subgroup of $G$ from Section 2.5; in particular, $G/G_{1}$ is isomorphic to a subgroup of $\operatorname{SL}_{2}(k)$. We have two cases: either $G_{1}$ is trivial, or $G_{1}\cong C_{3}$. In the first case it follows from the Dickson Theorem that $G$ is either a group from cases $(1-3)$, or is an extension of $C_{3}$ by a cyclic group of order dividing $80$. Since such an extension does not contain $C_{1}5$, its order divides $48$. As before, we get the following list of non-cyclic extensions: $D_{3}=C_{3}\rtimes C_{2}$, $Q_{12}=C_{3}\rtimes C_{4}$, and $C_{3}\rtimes C_{8}$. The group $D_{3}$ has no rigid representations according to Lemma 4.7.
In the second case, we have to consider extensions of $G_{1}\cong C_{3}$ by non-cyclic subgroups of $\operatorname{SL}_{2}(k)$ of order dividing $80$, i.e, $Q_{8}$, $Q_{16}$, $Q_{20}$, $Q_{40}$, or $Q_{80}$. Since $3$ is coprime to the orders of these groups, $H^{2}(G/G_{1},G_{1})$ is trivial; therefore, any extension is a semidirect product. The groups without non-small cyclic subgroups are $\operatorname{SL}_{2}({\mathbb{F}}_{3})\cong C_{3}\rtimes Q_{8}$, $C_{3}\times Q_{8}$, and $C_{3}\rtimes Q_{16}$.
Let $p=2$. The quotient group $\bar{G}=G/G_{1}$ is a subgroup of $\operatorname{SL}_{2}(k)$ without non-small cyclic subgroups; thus we have the following possibilities for $\bar{G}$:
,${}_{2},C_{3},C_{5},C_{6},C_{10},C_{2}\times C_{2},C_{6}\times C_{2},C_{10}\times C_{2},S_{3}=D_{3}=\operatorname{SL}_{2}({\mathbb{F}}_{2}),D_{5},A_{4}=C_{2}^{2}\rtimes C_{3},A_{5}=\operatorname{SL}_{2}({\mathbb{F}}_{4}).$
Assume first that $G_{1}$ is cyclic. If $\bar{G}$ is abelian, $D_{3}$, or $D_{5}$, then there is a normal subgroup $H$ of odd order such that $\bar{G}/H$ is a $2$-group. According to Lemma 4.6, any extension of $G_{1}$ by $\bar{G}$ is uniquely determined by the $2$-Sylow subgroup $Q$ of this extension. Since $Q$ is either cyclic or isomorphic to $Q_{8}$ or $Q_{16}$ we get the following groups without non-small cyclic subgroups:
  1. (1)
    $C_{4}$, $C_{8}$, or $Q_{16}$, if $\bar{G}=C_{2}$;
  2. (2)
    $C_{6}$, or $C_{12}$, if $\bar{G}=C_{3}$;
  3. (3)
    $C_{10}$, if $\bar{G}=C_{5}$;
  4. (4)
    $C_{12}$, if $\bar{G}=C_{6}$;
  5. (5)
    no groups, if $\bar{G}=C_{10}$;
  6. (6)
    $Q_{8}$, or $Q_{16}$, if $\bar{G}=C_{2}^{2}$;
  7. (7)
    $C_{3}\times Q_{8}$, if $\bar{G}=C_{6}\times C_{2}$;
  8. (8)
    no groups, if $\bar{G}=C_{10}\times C_{2}$;
  9. (9)
    $Q_{12}$, $Q_{24}$, or $C_{3}\rtimes C_{8}$, if $\bar{G}=D_{3}$,
  10. (10)
    $Q_{20}$, if $\bar{G}=D_{5}$.
If $\bar{G}$ is $A_{4}$ or $A_{5}$, then $G_{1}$ is $C_{2}$ or $C_{4}$. The only non-trivial extensions of $A_{4}$ and $A_{5}$ by $C_{2}$ are $\operatorname{SL}_{2}({\mathbb{F}}_{3})$ and $\operatorname{SL}_{2}({\mathbb{F}}_{5})$, and by $C_{4}$ are $C_{4}.A_{4}$ and $C_{4}.A_{5}$ respectfully. The groups $C_{4}.A_{4}$ and $C_{4}.A_{5}$ has no rigid representations according to Lemma 4.7.
If $G_{1}=Q_{16}$, then $\bar{G}$ is $C_{3}$ or $C_{5}$, and any such extension is trivial, because $\operatorname{Aut}(Q_{16})$ is a $2$-group; in this way we get two groups without non-small cyclic subgroups: $C_{3}\times Q_{16}$, and $C_{5}\times Q_{16}$. If $G_{1}=Q_{8}$, and $\bar{G}$ is $C_{3}$ or $C_{5}$, then either extension is trivial, or $G=Q_{8}\rtimes C_{3}\cong\operatorname{SL}_{2}({\mathbb{F}}_{3})$, because $\operatorname{Aut}(Q_{8})\cong S_{4}$.
Finally, assume that $G_{1}$ is $Q_{8}$, and the $2$-Sylow subgroup of $G$ is $Q_{16}$. If $\bar{G}$ is cyclic of even order, then any homomorphism from $\bar{G}$ to $\operatorname{Aut}(Q_{8})\cong S_{4}$ factors through $C_{2}$; therefore, either $G$ is $Q_{16}$, or contains a non-small cyclic subgroup. If $\bar{G}=C_{2}\times C_{2}$, then the order of the $2$-Sylow subgroup is greater than $16$. If $\bar{G}$ is $D_{3}$ or $D_{5}$, then any extension that contain $Q_{16}$ has a non-small cyclic subgroup.
We use the WEDDERGA package of GAP [We] to compute ${\mathbb{Q}}$-algebras for suitable representations without fixed points. ∎
5. Singularities of $A/G$
In this section we study resolutions of singularities of the quotients $A/G$. The next proposition is a generalization of the table from the Katsura’s paper [Ka87] for cyclic groups, and of results due to Fujiki [Fu] for $Q_{8}$, $Q_{12}$, and $\operatorname{SL}_{2}({\mathbb{F}}_{3})$.
Proposition 5.1.
Let $A$ be an abelian surface with a rigid and symplectic action of a group $G$ over $k$. Assume that $p$ does not divide the order of $G$. Then the singularities of the quotient surface $A/G$, and the rank of the Neron–Severi group of the minimal resolution of singularities $X$ are as follows.
$G$ Singularities $\operatorname{rk}\operatorname{NS}(X)$
$C_{2}$ $16A_{1}$ $\geq 17$
$C_{3}$ $9A_{2}$ $\geq 19$
$C_{4}$ $4A_{3}+6A_{1}$ $\geq 19$
$C_{5}$ $5A_{4}$ $22$
$C_{6}$ $A_{5}+4A_{2}+5A_{1}$ $\geq 19$
$C_{8}$ $2A_{7}+A_{3}+3A_{1}$ $22$
$C_{10}$ $A_{9}+2A_{4}+3A_{1}$ $22$
$C_{12}$ $A_{11}+A_{3}+2A_{2}+2A_{1}$ $22$
$Q_{8}$ $2D_{4}+3A_{3}+2A_{1}$ $\geq 20$
$4D_{4}+3A_{1}$ $\geq 20$
$Q_{12}$ $D_{5}+3A_{3}+2A_{2}+A_{1}$ $\geq 20$
$Q_{16}$ $2D_{6}+D_{4}+A_{3}+A_{1}$ $22$
$Q_{20}$ $D_{7}+A_{4}+3A_{3}$ $22$
$Q_{24}$ $D_{8}+D_{4}+2A_{3}+A_{2}$ $22$
$\operatorname{SL}_{2}({\mathbb{F}}_{3})$ $E_{6}+D_{4}+4A_{2}+A_{1}$ $\geq 20$
$\operatorname{ESL}_{2}({\mathbb{F}}_{3})$ $E_{7}+D_{6}+A_{3}+2A_{2}$ $22$
$\operatorname{SL}_{2}({\mathbb{F}}_{5})$ $E_{8}+D_{4}+A_{4}+2A_{2}$ $22$
Lemma 5.2.
Let $G$ be a finite group of order $n$ with a rigid action on an abelian variety $A$. Denote by $N(H)$ the number of points with the stabilizer isomorphic to $H$.
  1. (1)
    If $x$ is a fixed point of $G$, then there exists a prime $\ell$ such that $x\in A[\ell]({\bar{k}})$, and $n=\ell^{r}$ is a power of $\ell$.
  2. (2)
    If $G$ is cyclic and $n=\ell^{r}$, then the set of fixed points of $G$ is a subgroup of $A[\ell]({\bar{k}})$ of order $\ell^{r_{n}}$, where
    $r_{n}=\frac{2\dim A}{(\ell-1)\ell^{r-1}}$
  3. (3)
    If $H$ is a subgroup of $G$ such that $\ell$ does not divide the order of $H$, then the action of $H$ on $A[\ell]({\bar{k}})$ is free.
  4. (4)
    Let $G^{(\ell)}$ be the $\ell$-th Sylow subgroup of $G$. Suppose that there are $s_{\ell}$ congujacy classes of $\ell$-th Sylow subgroups. If $N_{\ell}$ is the number of points with stabilizer equal to $G^{(\ell)}$, then $N(G^{(\ell)})=s_{\ell}N_{\ell}$.
  5. (5)
    Let $G=C_{4}$, and let $A$ be a surface. Then $N(C_{4})=4$, and there are $6$ orbits of length $2$.
Proof.
Let $g\in G$ be an element of order $d$. Since the action is rigid, the Tate module $T_{\ell}(A)$ is a free module over the regular local ring ${\mathbb{Z}}_{\ell}[\zeta_{d}]$, where $\zeta_{d}$ acts as $g$. If $d=\ell$ is a prime, and $x$ is a fixed point of $G$, then $(\zeta_{\ell}-1)x=0$. Therefore, the norm
$N_{{\mathbb{Q}}(\zeta_{\ell})/{\mathbb{Q}}}(\zeta_{\ell}-1)=\ell$
also annihilates $x$. This proves $(1)$.
Let $G$ be cyclic of order $n=\ell^{r}$, and let $g\in G$ be a generator. Then $T_{\ell}(A)$ is a free module over ${\mathbb{Z}}_{\ell}[\zeta_{n}]$ of rank $r_{n}$. The set of fixed points of $g$ is isomorphic to
$T_{\ell}(A)/(g-1)T_{\ell}(A)\cong({\mathbb{Z}}_{\ell}[\zeta_{n}]/(\zeta_{n}-1){\mathbb{Z}}_{\ell}[\zeta_{n}])^{r_{n}}\cong{\mathbb{F}}_{\ell}^{r_{n}}.$
We proved $(2)$. Parts $(3)$ and $(4)$ are easy.
We now prove $(5)$. Clearly, $A[2]({\bar{k}})$ is fixed by $C_{2}\subset C_{4}$; therefore, according to $(2)$, we have $N(C_{4})=4$, and $12$ points form $6$ orbits of length $2$. ∎
Proof of Proposition 5.1.
We use the notation of Lemma 5.2. The table for cyclic groups can be easily computed using $(1),(2)$, and $(3)$ of Lemma 5.2.
Let $G=Q_{8}$, and let $H=C_{4}\subset G$ be a normal subgroup. According to Lemma 5.2.(1), we have to compute the orbits of $G$ on $A[2]({\bar{k}})$. The group $H$ has $4$ fixed points by Lemma 5.2.(5). Since the length of any non-trivial orbit is even, and the origin is fixed by $G$, either $N(G)=4$, or $N(G)=2$. Note that $Q_{8}$ contains $3$ cyclic subgroups of order $4$, and if $2$ such subgroups fix a point, then the point is fixed by $G$. In other words, either $N(C_{4})=0$, or $N(C_{4})=6$. In the first case, we have $N(C_{2})=12$, and in the second case $N(C_{2})=8$.
Let $G=Q_{16}$, and let $H=C_{8}\subset G$ be the only subgroup of order $8$. The action of $G/H$ on $A[2]({\bar{k}})/H$ is very simple, and we have $N(G)=N(C_{8})=2$, and $N(Q_{8})=2$. Moreover, since there are two conjugacy classes of $C_{4}$ in $G$, we have $N(C_{4})=4$, and $N(C_{2})=8$.
From the previous calculations we can easily compute the orbits in many cases using $(3)$ and $(4)$ of Lemma 5.2 and the following information:
  • $Q_{12}$ has $3$ cyclic $4$-Sylow subgroups, thus $N(G)=1$, $N(C_{4})=9$, $N(C_{3})=8$, and $N(C_{2})=6$;
  • $Q_{20}$ has $5$ cyclic $4$-Sylow subgroups, thus $N(G)=1$, $N(C_{5})=4$, and $N(C_{4})=15$;
  • $\operatorname{SL}_{2}({\mathbb{F}}_{3})$ has $4$ cyclic $3$-Sylow subgroups, and a normal subgroup isomorphic to $Q_{8}$, thus $N(G)=1$, $N(Q_{8})=3$, $N(C_{3})=32$, and $N(C_{2})=12$;
  • $\operatorname{SL}_{2}({\mathbb{F}}_{5})$ has $10$ cyclic $3$-Sylow subgroups, $6$ cyclic $5$-Sylow subgroups, and $5$ Sylow subgroups isomorphic to $Q_{8}$, thus $N(G)=1$, $N(Q_{8})=15$, $N(C_{5})=24$, and $N(C_{3})=80$.
Let $G=Q_{24}$. The action of $C_{2}\cong G/Q_{12}$ on $A[2]({\bar{k}})/Q_{12}$ clearly fixes the image of the origin and the image of a point with the stabilizer $C_{2}$; therefore, $N(G)=1$, and $N(C_{4})>1$. On the other hand, the action of $C_{2}\cong G/C_{12}$ on $A[2]({\bar{k}})/C_{12}$ fixes the image of a point with the stabilizer $C_{4}$; therefore, $N(Q_{8})=3$, and $N(C_{4})=12$.
Finally, let $G=\operatorname{ESL}({\mathbb{F}}_{3})$, and let $H=\operatorname{SL}_{2}({\mathbb{F}}_{3})$. Since the action of $C_{2}\cong G/H$ on $A[2]({\bar{k}})/H$ is trivial, $N(G)=1$, $N(Q_{16})=3$, and $N(C_{4})=12$. ∎
Remark 5.3.
There is a misprint for the cyclic group of order $8$ in [Ka87].
Corollary 5.4.
Let $A$ be an abelian surface over a perfect field $k$ of characteristic $p>2$. Assume that there exists a rigid and symplectic action on $A$ of a cyclic group $G$ of order $5,8$, or $12$. Then $A$ is supersingular.
If $p$ does not divide $n$, then ${\mathbb{F}}_{p^{2}}\subset k$.
Proof.
Let $X$ be the minimal resolution of singularities of $A/G$. If $p$ does not divide the order of $G$, then, according to the Katsura Theorem 1.3, and Proposition 5.1, $X$ is a supersingular $K3$ surface. By [Ka87, Lemma 4.4], $A$ is also supersingular.
If $p$ divides the order of $G$, and $A$ is mixed or ordinary, then either $A$ is simple and $\operatorname{End}^{\circ}(A)\cong{\mathbb{Q}}(\zeta_{n})$, or $A$ is isogenous to the square of an ordinary elliptic curve $B$ and
$\operatorname{End}^{\circ}(B)\subset{\mathbb{Q}}(\zeta_{n}).$
Furthermore, the slopes of $A$ are different; therefore, in the decomposition of $p{\mathbb{Z}}[\zeta_{n}]$ to the product of prime ideals there are at least two different primes. Since ${\mathbb{Z}}[\zeta_{n}]$ has only one prime ideal over $p$, when $p$ divides $n\in\{5,8,12\}$, the surface $A$ is supersingular.
Let ${\mathbb{H}}=\operatorname{End}^{\circ}(A)$. Assume that ${\mathbb{F}}_{p^{2}}\not\subset k$. Since ${\mathbb{Q}}[G]^{\mathrm{rig}}\cong{\mathbb{Q}}(\zeta_{n})$ is a subalgebra of ${\mathbb{H}}$, according to Theorems 2.7 and 2.8, we have the following possibilities:
  1. (1)
    ${\mathbb{H}}\cong{\mathbb{Q}}(\zeta_{n})$ is a field;
  2. (2)
    ${\mathbb{H}}\cong{\mathbb{H}}_{\infty}({\mathbb{Q}}(\sqrt{p}))$;
  3. (3)
    ${\mathbb{H}}\cong\operatorname{Mat}(2,{\mathbb{Q}}(\sqrt{-p}))$.
In the first case, according to Lemma 2.6, ${\mathbb{F}}_{p^{2}}\subset k$. A contradiction. In cases $(2)$ and $(3)$ the central simple algebra ${\mathbb{H}}$ of degree $8$ over ${\mathbb{Q}}$ contains ${\mathbb{Q}}(\zeta_{n})$ as a subfield of degree $4$ over ${\mathbb{Q}}$; therefore, the center of ${\mathbb{H}}$ is a subfield of ${\mathbb{Q}}(\zeta_{n})$, and ±$\sqrt{\pm p}\in{\mathbb{Q}}(\zeta_{n})$. This is impossible if $p$ does not divide $n$. ∎
Now we compute the Frobenius action on the graphs of exceptional curves on $X(A,G)$. Let $x\in A({\mathbb{F}}_{q^{r}})$ be a point of degree $r$, i.e., a morphism
$x:\operatorname{Spec}({\mathbb{F}}_{q^{r}})\to A.$
Let $S(x)$ be the stabilizer of $x$. For any $g\in S(x)$ there exists an element $\bar{g}\in\operatorname{Gal}({\mathbb{F}}_{q^{r}}/{\mathbb{F}}_{q})$ such that
$g\circ x=x\circ\bar{g}.$
We obtain a homomorphism
$\varphi_{x}:S(x)\to\operatorname{Gal}({\mathbb{F}}_{q^{r}}/{\mathbb{F}}_{q})$
given by the formula: $\varphi_{x}(g)=\bar{g}$.
Proposition 5.5.
  1. (1)
    Let $A$ be an abelian surface over ${\mathbb{F}}_{q}$ with a rigid and symplectic action of a cyclic group $G$ of order $n$ such that $(n,p)=1$. Let $x\in A({\mathbb{F}}_{q^{r}})$ be a point of degree $r$ fixed by $G$ such that $\varphi_{x}$ is surjective. Then the image of $x$ is an ${\mathbb{F}}_{q}$-point, the quotient singularity at $x$ is of type $A_{n/r-1}$, and the Frobenius action on the graph of exceptional curves is non-trivial if and only if $\zeta_{n/r}\not\in{\mathbb{F}}_{q^{r}}$.
  2. (2)
    Assume that $p>2$. Let $G=Q_{8}$, and let $A$ be an abelian surface over ${\mathbb{F}}_{q}$ with a rigid and symplectic action of $G$. Then the Frobenius action on the graph $D_{4}$ of the quotient singularity at the origin of $A$ is trivial.
Proof.
  1. (1)
    The algebra of the quotient singularity at $x$ is given by the formula
    $S=(\oplus_{s}(T_{x}^{*}(A))^{\otimes s})^{G}.$
    Let $k={\mathbb{F}}_{q^{r}}(\zeta_{n/r})$. We will compute $S_{k}=S\otimes_{{\mathbb{F}}_{q}}k$.
    Let $H=\operatorname{Gal}({\mathbb{F}}_{q^{r}}/{\mathbb{F}}_{q})$. By assumption, the homomorphism $\varphi_{x}:G\to H$ is surjective, and we can choose a generator $g\in G$ such that $\varphi_{x}(g)\in H$ is the Frobenius automorphism. There is an isomorphism of algebras
    ${\mathbb{F}}_{q^{r}}\otimes_{{\mathbb{F}}_{q}}k\cong\oplus_{h\in H}k_{h}$
    such that the Galois group $\operatorname{Gal}(k/{\mathbb{F}}_{q})$ acts through the natural surjection $\pi:\operatorname{Gal}(k/{\mathbb{F}}_{q})\to H$: namely, if $\lambda\in\operatorname{Gal}(k/{\mathbb{F}}_{q})$, then $\lambda(k_{h})=k_{\pi(\lambda)h}$. We obtain a corresponding decomposition for
    $T_{x}^{*}(A)\otimes_{{\mathbb{F}}_{q}}k\cong\oplus_{h\in H}V_{h},$
    where $V_{h}$ is a two-dimensional $k_{h}$-vector space with a semilinear action of $\ker\pi$. The action of $G$ is given by the formula $g(V_{h})=V_{\varphi_{x}(g)h}$ and the action of $\ker\varphi_{x}$ on $V_{h}$ is $k$-linear.
    Let $V=V_{e}$, where $e\in H$ is the trivial element. The natural injection
    $(\oplus_{s}V^{\otimes s})^{\ker\varphi_{x}}\to S_{k}$
    is a $\ker\pi$-equivariant isomorphism.
    Since $g^{r}\in\ker\varphi_{x}$, there exists a basis $u,v$ of $V$ such that $g^{r}v=\zeta v$, and $g^{r}u=\zeta^{-1}u$, where $\zeta\in k$ is a primitive root of unity of order $n/r$. Let $\lambda$ be a generator of $\ker\pi$. Then $\lambda(v)=\alpha u$ for some $\alpha\in k$ if and only if $\lambda(\zeta)\neq\zeta$, i.e, if and only if $\zeta\not\in{\mathbb{F}}_{q^{r}}$. In what follows we will assume that $\alpha=1$.
    It is straightforward to check that
    $x=v^{n/r},\;y=u^{n/r}\text{, and }z=uv$
    generate $\ker\varphi_{x}$-invariants of $\oplus_{s}V^{\otimes s}$ over $k$, and the equation of the singularity is $xy-z^{n/r}=0$. It follows that the Galois action on the exceptional graph becomes trivial over $k$. Moreover, the Galois action on the exceptional graph of the blow up is non-trivial if and only if $\lambda(x)=y$, i.e, if and only if $\zeta_{n/r}\not\in{\mathbb{F}}_{q^{r}}$.
  2. (2)
    Let $k={\mathbb{F}}_{q}(\zeta_{4})$. According to Lemma 6.7, there exists a basis $u,v$ of $T_{x}^{*}(A)\otimes_{{\mathbb{F}}_{q}}k$ such that $i$ and $j$ act as matrices
    $\left(\begin{array}[]{c c}\zeta_{4}&0\\ 0&-\zeta_{4}\\ \end{array}\right)\quad\text{ and }\quad\left(\begin{array}[]{c c}0&1\\ -1&0\\ \end{array}\right).$
    As before, the Galois group of $k$ over ${\mathbb{F}}_{q}$ interchanges $v$ and $\alpha u$ for some $\alpha\in{\mathbb{F}}_{q}^{*}$, and in what follows we may assume that $\alpha=1$. The $G$-invariants are generated by $x=v^{4}+u^{4}$, $y=v^{2}u^{2}$, and $z=vu(v^{4}-u^{4})$. Clearly, these functions are defined over ${\mathbb{F}}_{q}$. In these coordinates, the equation of the singularity is given by $z^{2}-y(x^{2}-4y^{2})=0$. The exceptional divisor of the blow up of the singular point is a projective line with three singular points of type $A_{1}$ defined over ${\mathbb{F}}_{q}$. Therefore, the Galois action on the resolution is trivial.
6. A refinement of the Katsura theorem.
In this section we use Corollary 3.9 to obtain a classification of finite groups that act on abelian varieties over a given finite field $k$ of characteristic $p$.
Proposition 6.1.
An algebra ${\mathbb{H}}$ from Theorem 4.3 admits a homomorphism to $M(2,{\mathbb{H}}_{p})$ if and only if $p$ satisfies the conditions in the table below.
[…]
${\mathbb{Q}}[\zeta_{4}]$ $p>0$ ${\mathbb{H}}_{2}$ $p>0$
${\mathbb{Q}}[\zeta_{3}]={\mathbb{Q}}[\zeta_{6}]$ $p>0$ ${\mathbb{H}}_{3}$ $p>0$
${\mathbb{Q}}[\zeta_{5}]={\mathbb{Q}}[\zeta_{10}]$ $p\not\equiv 1\bmod 5$ ${\mathbb{H}}_{\infty}({\mathbb{Q}}(\sqrt{5}))$ $p\not\equiv\pm 1\bmod 5$
${\mathbb{Q}}[\zeta_{8}]$ $p\not\equiv 1\bmod 8$ ${\mathbb{H}}_{\infty}({\mathbb{Q}}(\sqrt{2}))$ $p\not\equiv\pm 1\bmod 8$
${\mathbb{Q}}[\zeta_{12}]$ $p\not\equiv 1\bmod 12$ ${\mathbb{H}}_{\infty}({\mathbb{Q}}(\sqrt{3}))$ $p\not\equiv\pm 1\bmod 12$
Proof.
The first column of the table follows from Theorem 2.2 and the fact that ${\mathbb{Q}}(\zeta_{n})\otimes_{\mathbb{Q}}{\mathbb{H}}_{p}$ represents the trivial element of the Brauer group if and only if ${\mathbb{Q}}(\zeta_{n})$ does not split in $p$.
Let $r\in\{2,3\}$, but $r\neq p$. Put $K={\mathbb{Q}}(\sqrt{-rp})$. Clearly, both $K\otimes{\mathbb{H}}_{r}$, and $K\otimes{\mathbb{H}}_{p}$ are trivial. According to Theorem 2.2, there is a homomorphism $K\to{\mathbb{H}}_{p}$, and we have a sequence of homomorphisms of ${\mathbb{Q}}$-algebras:
${\mathbb{H}}_{r}\to{\mathbb{H}}_{r}\otimes K\cong M(2,K)\to M(2,{\mathbb{H}}_{p}).$
Finally, let $r\in\{2,3,5\}$. We prove that there is a homomorphism of ${\mathbb{H}}_{\infty}({\mathbb{Q}}(\sqrt{r}))$ to $M(2,{\mathbb{H}}_{p})$ if and only if $p$ does not split in ${\mathbb{Q}}(\sqrt{r})$. Indeed, there exists such a homomorphism if and only if there is a homomorphism in the opposite direction of the centralizers of these algebras in $M(8,{\mathbb{Q}})$, i.e., ${\mathbb{H}}_{p}\to{\mathbb{H}}_{\infty}({\mathbb{Q}}(\sqrt{r}))$. The result follows from Corollary 2.4. ∎
Remark 6.2.
There could be non-conjugate homomorphisms from ${\mathbb{Q}}^{\mathrm{rig}}[G]$ to $M(2,L)$, where $L$ is quadratic over ${\mathbb{Q}}$, but by the Skolem–Noether theorem, all homomorphisms from ${\mathbb{Q}}^{\mathrm{rig}}[G]$ to $M(2,{\mathbb{H}}_{p})$ are conjugate.
Theorem 6.3.
Assume that ${\mathbb{F}}_{p^{2}}$ is a subfield of a perfect field $k$, and the order of $G$ is prime to $p$ and is greater than $2$. Let $A$ be a supersingular abelian surface over $k$ with a rigid action of a finite group $G$. Then $G$ and $p$ satisfy the conditions in the column $I$. If the action is symplectic, then $G$ and $p$ satisfy the conditions in the column $II$.
If $p$ and $G$ satisfy a condition in the column $I$ of the table below, then there exists a supersingular abelian surface over ${\mathbb{F}}_{p^{2}}$ with a rigid action of a finite group $G$. If $p$ does not divide the order of $G$, then there exists a supersingular abelian surface over ${\mathbb{F}}_{p^{2}}$ with a rigid and symplectic action of a group $G$ if $p$ and $G$ satisfy the conditions in the column $II$.
$G$ $I$ $II$
$C_{3}$, $C_{6}$ $p>0$ $p>0$
$C_{4}$ $p>0$ $p>0$
$C_{8}$ $p\not\equiv 1\bmod 8$ $p\not\equiv\pm 1\bmod 8$
$C_{5}$ $C_{10}$ $p\not\equiv 1\bmod 5$ $p\not\equiv\pm 1\bmod 5$
$C_{12}$ $p\not\equiv 1\bmod 12$ $p\not\equiv\pm 1\bmod 12$,
$Q_{8}$ $p>0$ $p>0$
$Q_{12}$ $p>0$ $p>0$
$Q_{16}$ $p\not\equiv\pm 1\bmod 8$ $p\not\equiv\pm 1\bmod 8$
$Q_{20}$ $p\not\equiv\pm 1\bmod 5$ $p\not\equiv\pm 1\bmod 5$
$Q_{24}$ $p\not\equiv\pm 1\bmod 12$ $p\not\equiv\pm 1\bmod 12$
$\operatorname{SL}_{2}({\mathbb{F}}_{3})$ $p>0$ $p>0$
$\operatorname{ESL}_{2}({\mathbb{F}}_{3})$ $p\not\equiv\pm 1\bmod 8$ $p\not\equiv\pm 1\bmod 8$
$\operatorname{SL}_{2}({\mathbb{F}}_{5})$ $p\not\equiv\pm 1\bmod 5$ $p\not\equiv\pm 1\bmod 5$
First, we prove several lemmas in a slightly more general situation. Let $k$ be a perfect field, and let $A$ be an abelian surface over $k$ with a rigid action of a finite group $G$. If $p$ does not divide $n$, we denote by $\bar{\zeta}_{n}\in{\bar{k}}$ a primitive root of unity of degree $n$ such that $\bar{\zeta}_{n}$ lifts to $\zeta_{n}\in W({\bar{k}})$.
Lemma 6.4.
Assume that $k$ is finite. Let $g\in G$ be an element of order $p$. Then the image of $g$ in $\operatorname{GL}(T_{0}^{*}(A))\cong\operatorname{GL}_{2}(k)$ belongs to $\operatorname{SL}_{2}(k)$. In particular, the action of $g$ is symplectic.
Proof.
The index of $\operatorname{SL}_{2}(k)$ in $\operatorname{GL}_{2}(k)$ is prime to $p$. ∎
Lemma 6.5.
Let $g\in G$ be an element of order $n\in\{3,4,5,6,8,10,12\}$ prime to $p$. Assume that $\bar{\zeta}_{n}\not\in k$. Let ${\mathbb{F}}_{q}$ be the algebraic closure of ${\mathbb{F}}_{p}$ in $k$. Then the action on $T_{0}^{*}(A)$ is symplectic if and only if $q\equiv-1\bmod n$.
Proof.
The cotangent space is a one-dimensional vector space over $k(\bar{\zeta}_{n})$. The Galois group of $k(\bar{\zeta}_{n})$ over $k$ is of order $2$ and the action on $\bar{\zeta}_{n}$ is given by the formula:
$\bar{\zeta}_{n}\mapsto\bar{\zeta}_{n}^{q}.$
Therefore, the eigenvalues of a generator of $g$ are $\bar{\zeta}_{n}$ and $\bar{\zeta}_{n}^{q}$. In other words, the action is symplectic if and only if $q\equiv-1\bmod n$. ∎
Lemma 6.6.
Let $A$ be a non-ordinary surface, and let $g\in G$ be an element of order $n\in\{3,4,5,6,8,10,12\}$ prime to $p$. Assume that $\bar{\zeta}=\bar{\zeta}_{n}\in k$.
  1. (1)
    If $p^{2}\equiv 1\mod n$, then there exists a basis $v_{1},v_{2},u_{1},u_{2}$ of $M(A)\otimes{\mathbb{Q}}$ such that
    ()( ∗ ) $g(v_{1})=\zeta v_{1},\quad g(v_{2})=\zeta^{p}v_{2},\quad g(u_{1})=\zeta^{\prime}u_{1},\quad g(u_{2})=\zeta^{\prime p}u_{2},$
    where $\zeta^{\prime}$ is a prime to $n$ power of $\zeta$. If the action on $T_{0}^{*}(A)$ is symplectic, then there exists a basis such that ()(*)( ∗ ) holds with $\zeta^{\prime}=\zeta^{-1}$.
  2. (2)
    If $n\in\{5,8,10,12\}$, then $p\not\equiv\pm 1\bmod n$.
  3. (3)
    If $n\in\{3,4,6\}$, or $n\in\{8,12\}$, and $p\not\equiv\pm 1\bmod n$, then we can choose a basis of $M(A)\otimes{\mathbb{Q}}$ such that ()(*)( ∗ ) holds with $\zeta^{\prime}=\zeta^{-1}$, and
    $F(v_{1})=v_{2},\text{ and }F(u_{1})=u_{2}.$
Proof.
There exists a basis $\bar{v}_{1},\bar{u}_{1}$ of the cotangent space such that
$g(\bar{v}_{1})=\bar{\zeta}\bar{v}_{1}\text{, and }g(\bar{u}_{1})=\bar{\zeta}^{\prime}\bar{u}_{1},$
where $\bar{\zeta}^{\prime}\in k$ is a power of $\bar{\zeta}$. We can lift this basis to elements $v_{1}$ and $u_{1}$ of $M(A)$ such that
$g(v_{1})=\zeta v_{1}\text{, and }g(u_{1})=\zeta^{\prime}u_{1},$
where $\zeta^{\prime}\in W(k)$ lifts $\bar{\zeta}^{\prime}$. Put $v_{2}=F(v_{1})$, and $u_{2}=F(u_{1})$. Then
$g(v_{2})=\zeta^{p}v_{2}\text{, and }g(u_{2})=(\zeta^{\prime})^{p}u_{2}.$
Assume that $p^{2}\equiv 1\mod n$. Then the equality $\zeta^{p}=\zeta^{\prime}$ is equivalent to $(\zeta^{\prime})^{p}=\zeta$, because $(\zeta^{\prime})^{p}=\zeta^{p^{2}}=\zeta$. It follows that, if $\zeta^{p}=\zeta^{\prime}$, then $v_{1},u_{1}$ generate a Diedonné submodule of $M(A)$ of slope $1$. By duality, slopes of $A$ are 00 and $1$, i.e., $A$ is ordinary. A contradiction. Therefore, $\zeta^{p}\neq\zeta^{\prime}$, and $v_{1},v_{2},u_{1},u_{2}$ is a basis of $M(A)\otimes{\mathbb{Q}}$. Clearly, the action of $g$ on $T_{0}^{*}(A)$ is symplectic if and only if $\zeta^{\prime}=\zeta^{-1}$. Part $(1)$ is proved.
We now prove $(2)$. If $n\in\{5,8,10,12\}$, and $v_{1},v_{2},u_{1},u_{2}$ is a basis of $M(A)\otimes{\mathbb{Q}}$, then the eigenvalues of $g$ are different by rigidity, and $p\not\equiv\pm 1\bmod n$. Suppose that $v_{1},v_{2},u_{1},u_{2}$ is not a basis of $M(A)\otimes{\mathbb{Q}}$. According to part $(1)$, we have $p^{2}\not\equiv 1\mod n$, i.e., $n\in\{5,10\}$, and $p\not\equiv\pm 1\mod 5$. We proved that if the action of $g$ on $T_{0}^{*}(A)$ is symplectic, then $p\not\equiv\pm 1\bmod n$.
Now, we construct a basis of $M(A)\otimes{\mathbb{Q}}$ that satisfy conditions of $(3)$; we call such a basis symplectic. Our assumptions imply that $p^{2}\equiv 1\mod n$. We already proved that in this case a basis $v_{1},v_{2},u_{1},u_{2}$ with property ()(*)( ∗ ) exists.
Assume that $n\in\{3,4,6\}$. According to Lemma 3.6, either $\zeta^{\prime}=\zeta^{-1}$ and the basis is symplectic, or $\zeta^{\prime}=\zeta^{-p}$. In the second case, the basis $v_{1},v_{2},u_{2},F(u_{2})$ is symplectic.
Assume that $n\in\{5,8,10,12\}$, and $p\not\equiv\pm 1\bmod n$. According to Lemma 3.6, all eigenvalues of $g$ are different; therefore, either $\zeta^{\prime}=\zeta^{-1}$ and the basis is symplectic, or $\zeta^{\prime}=\zeta^{-p}$. In the second case the basis $v_{1},v_{2},u_{2},F(u_{2})$ is symplectic. The lemma is proved. ∎
Lemma 6.7.
Let $G=Q_{4n}$, and let $g\in G$ be a element of order $2n$. Assume that $2n$ is prime to $p$. Then over $k(\bar{\zeta}_{2n})$ there is a basis of $T_{0}^{*}(A)$ such that generators of $G$ act as matrices
$\left(\begin{array}[]{c c}\bar{\zeta}_{2n}&0\\ 0&\bar{\zeta}_{2n}^{-1}\\ \end{array}\right)\quad\text{ and }\quad\left(\begin{array}[]{c c}0&1\\ -1&0\\ \end{array}\right).$
In particular, the action of $G$ is symplectic.
Proof.
There exists an element $j\in G$ of order $4$ such that $jgj^{-1}=g^{-1}$. Therefore, $\bar{\zeta}_{2n}^{-1}$ is also an eigenvalue of $g$, and there exists a basis $v,u\in T^{*}(A)\otimes k(\bar{\zeta}_{2n})$ such that the matrix of $g$ is diagonal. It is straightforward to show that $j(v)=\gamma u$ for some $\gamma\in k(\bar{\zeta}_{2n})$, and $j(u)=-\gamma^{-1}v$. Therefore, the basis $v,\gamma u$ is a desired basis. ∎
The following lemma is well known.
Lemma 6.8.
If $p>3$, then there exists a basis of $T_{0}^{*}(A)$ such that $\operatorname{SL}_{2}({\mathbb{F}}_{3})$ is generated by $Q_{8}$ and the matrix
$\frac{1}{2}\left(\begin{array}[]{c c}1+\zeta_{4}&1+\zeta_{4}\\ -1+\zeta_{4}&1-\zeta_{4}\\ \end{array}\right);$
and $\operatorname{ESL}_{2}({\mathbb{F}}_{3})$ is generated by $\operatorname{SL}_{2}({\mathbb{F}}_{3})$ and the matrix
$\left(\begin{array}[]{c c}\zeta_{8}&0\\ 0&\zeta_{8}^{-1}\\ \end{array}\right).$
If $p>5$, then there exists a basis of $T_{0}^{*}(A)$ such that $\operatorname{SL}_{2}({\mathbb{F}}_{5})$ is generated by $\operatorname{SL}_{2}({\mathbb{F}}_{3})$ and the matrix
$\left(\begin{array}[]{c c}\zeta_{5}&0\\ 0&\zeta_{5}^{-1}\\ \end{array}\right).$
These matrices are symplectic; therefore, a symplectic basis for $Q_{8}$ is a symplectic basis for $\operatorname{SL}_{2}({\mathbb{F}}_{3})$, $\operatorname{ESL}_{2}({\mathbb{F}}_{3})$, or $\operatorname{SL}_{2}({\mathbb{F}}_{5})$ as well. ∎
Proof of theorem 6.3..
If there exists a supersingular abelian surface with a rigid action of a group $G$, then, according to Corollary 3.10, there exists a supersingular abelian surface $B$ over ${\mathbb{F}}_{p^{2}}$ with a rigid action of $G$. It follows that $G$ and $p$ satisfy the conditions in column $I$ by Theorem 4.3, and Proposition 6.1. Let $A$ be an abelian surface over $k$ with a rigid and symplectic action of $G$. Then the conditions in column $II$ follow from Lemmata 6.6 and 6.5.
Let us prove the existence part. According to Theorem 2.7, there exists a supersingular elliptic curve $E$ over ${\mathbb{F}}_{p^{2}}$ such that $\operatorname{End}^{\circ}(E)\cong{\mathbb{H}}_{p}$; it follows that $\operatorname{End}^{\circ}(E^{2})\cong M(2,{\mathbb{H}}_{p}).$ By Corollary 3.9, there exists a supersingular abelian surface with an action of a group $G$ if and only if there exists a homomorphism ${\mathbb{Q}}[G]^{\mathrm{rig}}\to M(2,{\mathbb{H}}_{p})$. According to Theorem 4.3, and Proposition 6.1, the conditions in the first column of the table imply that such a homomorphism exists.
Let $A=E^{2}$, and let ${\mathbb{H}}={\mathbb{Q}}[G]^{\mathrm{rig}}$. We are going to prove that under the conditions of column $II$ there exists a submodule $M$ of $M(A)\otimes{\mathbb{Q}}$ such that the action on $M/FM$ is symplectic and $M\otimes{\mathbb{Q}}\cong M(A)\otimes{\mathbb{Q}}$; we will call such a submodule symplectic. Assume that a symplectic submodule exists. Then, according to Lemma 2.9, there exist an abelian variety $B$ and a $p$-isogeny $\varphi:A\to B$ such that $M(B)\cong M$; in particular, for all $\ell\neq p$ we have $T_{\ell}(A)\cong T_{\ell}(B)$. Therefore, the natural homomorphism $G\to\operatorname{End}^{\circ}(B)$ factors through $\operatorname{End}(B)$, and $G$ acts on $B$. By construction, the isogeny $\varphi$ is $G$–equivariant, and the action of $G$ on $B$ is rigid and symplectic.
Now, we construct a symplectic submodule. Firstly, assume that $G$ is cyclic of order $n$. Let ${\mathbb{H}}={\mathbb{Q}}(\zeta_{n})$ be a quadratic field extension of ${\mathbb{Q}}$. If $\bar{\zeta}_{n}\not\in k$, then the action is always symplectic according to Lemma 6.5. If $\bar{\zeta}_{n}\in k$, then, according to Lemma 6.6, we can choose a symplectic submodule $M$ of $M(A)\otimes{\mathbb{Q}}$.
Assume that ${\mathbb{H}}={\mathbb{Q}}(\zeta_{n})$ is a quartic field extension of ${\mathbb{Q}}$, i.e., $n\in\{5,8,10,12\}$. Then either $n$ divides $p^{2}-1$, or ${\mathbb{H}}\cong{\mathbb{Q}}(\zeta_{5})$. In the first case, Lemma 6.6 implies that there exists a symplectic submodule $M$ in $M(A)$. If ${\mathbb{H}}\cong{\mathbb{Q}}(\zeta_{5})$, then according to Lemma 6.5 the action is symplectic if and only if $p^{2}\equiv-1\bmod 5$, i.e, $p\not\equiv\pm 1\bmod 5$.
The remaining cases follow from Lemmata 6.7 and 6.8. ∎
Proposition 6.9.
Let ${\mathbb{H}}$ be a central simple algebra over ${\mathbb{Q}}$, and let $A$ be an ordinary abelian variety over $k$ with a rigid action of a finite group $G$ such that the action factors through ${\mathbb{H}}:$
${\mathbb{Q}}[G]^{\mathrm{rig}}\to{\mathbb{H}}\to\operatorname{End}^{\circ}(A).$
Then the characteristic polynomial of the action of any $g\in G$ on $T_{0}^{*}(A)$ is congruent modulo $p$ to a power of the cyclotomic polynomial $\Phi_{n}$, where $n$ is the order of $g$.
Proof.
Let $V$ be the irreducible representation of ${\mathbb{H}}$ over ${\mathbb{Q}}$. The induced representation of $G$ on $V$ is without fixed points, hence the characteristic polynomial of any element $g\in G$ of order $n$ is equal to some power of the cyclotomic polynomial $\Phi_{n}^{r_{g}}$. According to Proposition 2.12, $M(A)\cong M\oplus M^{*}$ is a direct sum of Diedonné modules. The action of $G$ on $M$ comes from the composition
${\mathbb{Q}}[G]^{\mathrm{rig}}\to{\mathbb{H}}\otimes{\mathbb{Q}}_{p}\to\operatorname{End}^{\circ}(A)\otimes{\mathbb{Q}}_{p}\to\operatorname{End}(M\otimes{\mathbb{Q}}_{p}).$
It follows that $M\otimes{\mathbb{Q}}_{p}$ is isomorphic to $(V\otimes{\mathbb{Q}}_{p})^{r_{M}}$ as a ptensor-${\mathbb{H}}\otimes{\mathbb{Q}}_{p}$-module, and the characteristic polynomial of the action of $g$ on $M\otimes{\mathbb{Q}}_{p}$ is equal to $\Phi_{n}^{r_{g}r_{M}}$. According to Proposition 2.12, we have the isomorphism $T^{*}(A)\cong M/pM$. The proposition is proved. ∎
Example 6.10.
Let $V$ be a ${\mathbb{Q}}$-representation of $G$ without fixed points, and let ${\mathbb{H}}=M(d,{\mathbb{Q}})$, where $d=\dim V$. For any abelian variety $B$ over $k$ we get homomorphisms
${\mathbb{Q}}[G]^{\mathrm{rig}}\to M(d,{\mathbb{Q}})\to M(d,{\mathbb{Q}})\otimes\operatorname{End}^{\circ}(B)\cong\operatorname{End}^{\circ}(B^{d}),$
According to Corollary 3.9, there exists an abelian variety $B_{V}$ of dimension $d\dim B$ with a rigid action of $G$.
Theorem 6.11.
There exists an abelian surface over ${\mathbb{F}}_{p}$ with a rigid and symplectic action of $G$ in the following cases:
  1. (1)
    $G$ is a cyclic group of order $2,3,4$, or $6$;
  2. (2)
    $p\neq 2$, and $G$ is $Q_{8}$, or $\operatorname{SL}_{2}({\mathbb{F}}_{3})$;
  3. (3)
    $p>3$, and $G\cong Q_{12}$.
Proof.
Let $E$ be an ordinary elliptic curve over ${\mathbb{F}}_{p}$, and let $G=C_{n}$, where $n\in\{3,4,6\}$. The representation of $G$ on $V={\mathbb{Q}}^{\mathrm{rig}}[G]$ is a two-dimensional representation without fixed points. The abelian variety $E_{V}$ constructed in Example 6.10 is a surface, and, by Proposition 6.9, the characteristic polynomial of the action of a generator on $T^{*}_{0}(A_{V})$ is equal to $\Phi_{n}$. According to Lemma 6.4, the action is rigid and symplectic. This proves part $(1)$.
If $G$ is not cyclic, we will construct an ordinary elliptic curve $E$ such that there exists a homomorphism from ${\mathbb{Q}}^{\mathrm{rig}}[G]$ to $\operatorname{End}^{\circ}(E^{2})\cong M(2,L)$, where $L=\operatorname{End}^{\circ}(E)$. Then, according to Corollary 3.9, there exists an abelian surface $A$ isogenous to $E^{2}$ with a rigid action of $G$. If $G$ is $Q_{8},Q_{12}$, or $\operatorname{SL}_{2}({\mathbb{F}}_{3})$, then we use Proposition 6.9 with ${\mathbb{H}}={\mathbb{Q}}[G]^{\mathrm{rig}}$ and find that the action on $A$ is symplectic.
Let $G$ be $Q_{8}$ or $\operatorname{SL}_{2}({\mathbb{F}}_{3})$. According to Corollary 2.4, if $L$ is not split in $2$, then there exists a homomorphism from ${\mathbb{Q}}^{\mathrm{rig}}[G]\cong{\mathbb{H}}_{2}$ to $M(2,L)$. We are going to apply Theorem 2.7 and find an elliptic curve $E$ such that its endomorphism algebra $L$ is not split in $2$. If $p\equiv 1\bmod 4$, then there exists an elliptic curve $E$ with the Weil polynomial $t^{2}-t+p$; the discriminant of this polynomial is $1-4p\equiv 5\bmod 8$. If $p\equiv 3\bmod 4$, then there exists an elliptic curve $E$ with the Weil polynomial $t^{2}-4t+p$; the discriminant of this polynomial is $16-4p\equiv 5\bmod 8$. In both cases $L$ is inert in $2$, if $p>2$.
If $G=Q_{12}$, then ${\mathbb{Q}}^{\mathrm{rig}}[G]={\mathbb{H}}_{3}$. As before, we need an ordinary elliptic curve over ${\mathbb{F}}_{p}$ such that its endomorphism algebra $L$ is not split in $3$. If $p\equiv 1\bmod 3$, then there exists an elliptic curve $E$ with the Weil polynomial $t^{2}-3t+p$; the discriminant of this polynomial is $9-4p\equiv 2\bmod 3$. If $p\equiv 2\bmod 3$, then there exists an elliptic curve $E$ with the Weil polynomial $t^{2}-t+p$; the discriminant of this polynomial is $1-4p\equiv 2\bmod 3$. In both cases $L$ is inert in $3$, if $p>3$. ∎
Proof of Theorem 1.5..
The existence part follows from Theorems 6.3 and 6.11. On the other hand, according to Theorem 1.3, the action of $G$ on $A$ is rigid and symplectic. Since $p$ does not divide the order of $G$, it follows from Theorem 4.3 that the group $G$ belongs to the Katsura list from Theorem 1.4.
If $G$ contains a cyclic subgroup of order $n\in\{5,8,12\}$, then, according to Corollary 5.4, the abelian surface $A$ is supersingular, and ${\mathbb{F}}_{p^{2}}\subset k$. In this situation, according to Theorem 6.3, we have $p\not\equiv\pm 1\bmod n$. ∎
7. Traces of Frobenius
In this section, we use our methods to compute traces of Frobenius and zeta functions of generalized Kummer surfaces.
7.1. Zeta functions of abelian varieties and generalized Kummer surfaces
Let $k={\mathbb{F}_{q}}$ be a finite field of order $q$, and let $X$ be an algebraic variety over $k$. Denote by $N_{r}$ the number of points on $\bar{X}$ defined over ${\mathbb{F}}_{q^{r}}$. Then the zeta function of $X$ is the formal power series
$Z_{X}(t)=\exp(\sum_{r=1}^{\infty}\frac{N_{r}t^{r}}{r}).$
The zeta function of an abelian variety $A$ with the Weil polynomial $f_{A}=\prod_{j}(t-\pi_{j})$ can be computed as follows:
$Z_{A}(t)=\prod_{i=0}^{2\dim A}P_{i}(A,t)^{(-1)^{i+1}},$
where
$P_{i}(A,t)=\prod_{j_{1}<\dots<j_{i}}(1-t\pi_{j_{1}}\dots\pi_{j_{i}}).$
The zeta function of a Shioda supersingular $K3$ surface is given by the formula
(7.1) $Z_{X}(t)=\frac{1}{(1-t)P_{2}(qt)(1-q^{2}t)},$
where
$P_{2}(t)=\det(1-tF|\operatorname{NS}({\overline{X}})\otimes{\mathbb{Q}})=\prod_{r}\Phi_{r}(t)^{\lambda_{r}}$
is the characteristic polynomial of Frobenius automorphism on $\operatorname{NS}({\overline{X}})$. Since the roots of this polynomial are roots of unity, we can write it as a product of cyclotomic polynomials $\Phi_{r}$. Therefore, the zeta function is uniquely determined by $P_{2}$, and this polynomial is determined by the numbers $\lambda_{r}$. We denote the zeta function by $(1^{\lambda_{1}},\dots,r^{\lambda_{r}})$.
It follows from the definitions that the number of points on a supersingular $K3$ surface $X$ over ${\mathbb{F}}_{q}$ is equal to
$1+q\operatorname{Tr}_{X}+q^{2},$
where $\operatorname{Tr}_{X}$ is the trace of the Frobenius action on $\operatorname{NS}({\overline{X}})$.
7.2. Traces of Frobenius over ${\mathbb{F}}_{p^{2r}}$.
The following lemma is well known.
Lemma 7.1.
Let $X=X(A,G)$ be a generalized Kummer surface. Then
$\operatorname{NS}({\overline{X}})\otimes{\mathbb{Q}}\cong\oplus_{E}{\mathbb{Q}}E\oplus\operatorname{NS}(\overline{A})^{G}\otimes{\mathbb{Q}},$
where the sum is over the exceptional lines $E$ of the resolution of singularities of $A/G$.
First, we compute the zeta function of the quotient $A/G$.
Lemma 7.2.
Let $A$ be an abelian surface over ${\mathbb{F}}_{q}$ with the Weil polynomial $f(t)=(t^{2}+\varepsilon q)^{2}$ with a rigid action of a group $G$, where $\varepsilon=\pm 1$. Denote the characteristic polynomial of the Frobenius action on $H^{2}(\bar{A},{\mathbb{Q}}_{2})^{G}$ by $h(t)$.
  • If $G$ is cyclic of order $3,4$, or $6$, then $h(t)=(t-q)^{2}(t+q)^{2}$.
  • If $G=Q_{8}$, $Q_{12}$, or $\operatorname{SL}_{2}({\mathbb{F}}_{3})$, then $h(t)=(t+\varepsilon q)^{2}(t-\varepsilon q)$.
Proof.
Let $V=H^{1}(\bar{A},{\mathbb{Q}}_{2})\otimes{\mathbb{Q}}(\sqrt{-\varepsilon q})$. Then $V=V_{1}\oplus V_{2}$, where $F$ acts as $(-1)^{s}\sqrt{-\varepsilon q}$ on $V_{s}$, and $s\in\{1,2\}$. Since the action of $G$ commute with the Frobenius action, both $V_{1}$ and $V_{2}$ are $G$-invariant, and
$\wedge^{2}(V)^{G}=\wedge^{2}V_{1}\oplus\wedge^{2}V_{2}\oplus(V_{1}\otimes V_{2})^{G}.$
If $G$ is a cyclic group, then $\dim(V_{1}\otimes V_{2})^{G}=2$. This proves (1). In case (2) it is not hard to check that $\dim(V_{1}\otimes V_{2})^{G}=1$. ∎
Proposition 7.3.
Let $k={\mathbb{F}}_{q}$, where $q\equiv 1\bmod 4$. Then there exist a supersingular generalized Kummer surfaces $K(A,C_{4})$ over $k$ with the zeta functions $(1^{15},2^{7})$, and $(1^{20},2^{2})$. In both cases, the Weil polynomial of $A$ is equal to $(t^{2}-q)^{2}$.
Proof.
Let $A^{\prime}$ be an abelian surface with the Weil polynomial $(t^{2}-q)^{2}$. According to Lemma 2.2, there exists a homomorphism from ${\mathbb{Q}}[\zeta_{4}]$ to
$\operatorname{End}^{\circ}(A)\cong{\mathbb{H}}_{\infty}({\mathbb{Q}}(\sqrt{p}).$
According to Corollary 3.9 and Lemma 6.5, there exists an isogeny from $A^{\prime}$ to an abelian surface $A$ with a rigid and symplectic action of $C_{4}$. Denote by $g\in\operatorname{End}(A)$ the image of a generator of $C_{4}$.
Since $F$ and $g$ commute, the vector space $V_{2}(A)=V_{1}\oplus V_{2}$ is a sum of $2$-dimensional $F$-invariant subspaces such that $g(V_{1})=V_{2}$. According to [Ry14, Proposition 3.5], there exists an $F$-invariant ${\mathbb{Z}}_{2}$-submodule $T^{\prime}\subset V_{1}$ such that the action of $F$ on $T^{\prime}/2T^{\prime}$ is non-trivial, but the action of $F^{2}$ is trivial. Moreover, if $q\equiv 1\bmod 4$, then there exists an $F$-invariant ${\mathbb{Z}}_{2}$-submodule $T\subset V_{1}$ such that $F$ acts on $T/2T\cong({\mathbb{Z}}/2{\mathbb{Z}})^{2}$ as identity.
By Lemma 2.5, there exists an abelian surface $B$ with a $2$-isogeny $A\to B$ and such that $T\oplus gT\cong T_{2}(B)$. In particular, there exists a rigid and symplectic action of $C_{4}$ on $B$. We have $B[2](k)\cong({\mathbb{Z}}/2{\mathbb{Z}})^{16}$ and, according to Proposition 5.1, the singularities of $A/G$ are $4A_{3}+6A_{1}$; therefore, according to Lemma 7.2 and Proposition 5.5, the zeta function of the generalized Kummer surface $K(B,C_{4})$ is $(1^{20},2^{2})$.
In the same way, there exists an abelian surface $B^{\prime}$ with a rigid and symplectic action of $G$ such that $T^{\prime}\oplus gT^{\prime}\cong T_{2}(B^{\prime})$. A straightforward computation shows that on $B^{\prime}/G$ there are exactly two $F$-invariant singular points of type $A_{3}$ and two $F$-invariant singular points of type $A_{1}$. Moreover, since the action of $F^{2}$ is trivial on $B^{\prime}[2]({\bar{k}})$, other singular points are of degree $2$ over $k$. According to Proposition 5.5, there are $8$ $F$-invariant exceptional lines, and $5$ pairs of lines such that Frobenius acts non-trivially. It follows that, the zeta function of $K(B^{\prime},C_{4})$ is $(1^{15},2^{7})$. ∎
Theorem 7.4.
Assume that $k={\mathbb{F}}_{q}$, where $q=p^{2r}$, and $p>2$. There exists a generalized Kummer surface $X$ over $k$ corresponding to an abelian surface $A$ with the following parameters.
$\operatorname{Tr}_{X}$ $Z_{X}(t)$ $G$ $p$ $f_{A}(t)$
$22$ $1^{22}$ $C_{2}$ $p>2$ $(t\pm\sqrt{q})^{4}$
$18$ $1^{20},2^{2}$ $C_{4}$ $p>2$ $(t^{2}-q)^{2}$
$14$ $1^{18},2^{4}$ $C_{2}$ $p>2$ $(t^{2}-q)^{2}$
$10$ $1^{14},2^{4},4^{4}$ $C_{2}$ $p>2$ $(t^{2}+q)(t\pm\sqrt{q})^{2}$
$8$ $1^{15},2^{7}$ $C_{4}$ $p>2$ $(t^{2}-q)^{2}$
$6$ $1^{14},2^{8}$ $C_{2}$ $p>2$ $(t^{2}\pm q)^{2}$
$4$ $21^{10},3^{1}$ 2 $C_{2}$ $p>2$ $(t^{2}\pm\sqrt{q}t+q)^{2}$
$2$ $1^{12},2^{10}$ $C_{2}$ $p>2$ $(t^{2}-q)^{2}$
00 $1^{6},2^{4},3^{8},6^{4}$ $C_{2}$ $p\not\equiv 1\bmod 12$ $t^{4}-qt^{2}+q^{2}$
Proof.
The results for $G=C_{2}$ follow from Theorem 2.8, and [Ry14, Theorem 7.1], and for $G=C_{4}$ from Proposition 7.3. ∎
7.3. Traces of Frobenius over ${\mathbb{F}}_{p^{2r+1}}$.
In this section $k={\mathbb{F}}_{p^{2r+1}}$ is an odd extension of ${\mathbb{F}}_{p}$. We start this section with a nonexistence result over $k$.
Theorem 7.5.
[Ar] Suppose $p\neq 2$. If a supersingular $K3$ surface $X$ is defined over ${\mathbb{F}}_{q}$ and Frobenius acts trivially on $\operatorname{NS}({\overline{X}})$, then $p^{2}$ divides $q$.
Corollary 7.6.
If $X$ is a Shioda supersingular $K3$ surface over $k={\mathbb{F}}_{p^{2r+1}}$, then $\operatorname{Tr}_{X}\neq 22$.
Corollary 7.7.
If $q$ is an odd power of $p$, then in the product $P_{2}(t)=\prod_{r}\Phi_{r}(t)^{\lambda_{r}}$ there exists an even $r$ such that $\lambda_{r}>0$.
Proof.
Suppose that for all even $r$ we have $\lambda_{r}=0$. There exists an odd $m$ such that $F^{m}$ acts trivially on $\operatorname{NS}({\overline{X}})$. We get a contradiction to the Artin Theorem for the surface $X\otimes_{{\mathbb{F}}_{q}}{\mathbb{F}}_{q^{m}}$. ∎
Corollary 7.8.
Let $X$ is a supersingular $K3$ surface over $k$ with $\operatorname{Tr}_{X}=21$. Then the zeta function of $X$ is equal to $(1^{20},6^{2})$.
Question 7.9.
Is it true that $\operatorname{Tr}_{X}\neq 21$, and $\operatorname{Tr}_{X}\neq 19$ for any supersingular $K3$ surface over ${\mathbb{F}}_{q}$, where $q$ is an odd power of $p$?
We now prove an analog of Theorem 6.3 over a finite field of odd degree.
Theorem 7.10.
Let $k={\mathbb{F}}_{p^{2r+1}}$. Assume that the order of $G$ is greater than $2$ and is not divisible by $p$. There exists a supersingular abelian surface with the Weil polynomial $f$ and a rigid action of a group $G$ if and only if $p$, $f$, and $G$ satisfy one of the following conditions.
  1. (1)
    If $f(t)=t^{4}+q^{2}$, then $G=C_{4}$.
  2. (2)
    If $f(t)=t^{4}\pm qt^{2}+q^{2}$, then $G=C_{3}$, or $G=C_{6}$.
  3. (3)
    If $f(t)=(t^{2}-q)^{2}$, then $G$ and $p$ satisfy the following conditions:
    $G$ $p$
    $C_{3}$, $C_{4}$, $C_{6}$ $p>2$
    $Q_{8}$ $p\not\equiv 1\bmod 8$
    $Q_{12}$ $p\not\equiv 2\bmod 3$
    $\operatorname{SL}_{2}({\mathbb{F}}_{3})$ $p\not\equiv 1\bmod 8$
  4. (4)
    If $f(t)=(t^{2}+q)^{2}$, then $G$ and $p$ satisfy the following conditions:
    $G$ $p$
    $C_{3}$, $C_{4}$, $C_{6}$ $p>2$
    $Q_{8}$ $p\not\equiv-1\bmod 8$
    $Q_{12}$ $p\not\equiv 1\bmod 3$
    $\operatorname{SL}_{2}({\mathbb{F}}_{3})$ $p\not\equiv-1\bmod 8$
  5. (5)
    $f(t)=(t^{2}\pm 3^{r}+q)^{2}$, and $p=3$; in this case, $G$ is isomorphic to $C_{4}$, $C_{8}$, or $Q_{8}$;
  6. (6)
    $f(t)=(t^{2}\pm 2^{r}+q)^{2}$, and $p=2$; in this case, $G\cong C_{3}$.
In all these cases, there exists an abelian surface with a rigid and symplectic action of the group $G$.
Proof.
Suppose that $G$ acts on a supersingular surface $A$ over $k$. The surface $A$ is either simple or isogenous to a product of two elliptic curves, say $E_{1}$ and $E_{2}$. If $E_{1}$ and $E_{2}$ are not isogenous, and the action of $G$ is rigid, there are homomorphisms
${\mathbb{Q}}[G]^{\mathrm{rig}}\to\operatorname{End}^{\circ}(E_{i})$
for $i\in\{1,2\}$. According to Theorem 2.7, there are three possibilities for $\operatorname{End}^{\circ}(E_{i})$:
  • ${\mathbb{Q}}(\sqrt{-p})$;
  • ${\mathbb{Q}}(\zeta_{3})$, and $p=3$;
  • ${\mathbb{Q}}(\zeta_{4})$, and $p=2$.
Therefore, $G$ is abelian, and according to Corollary 3.13, $G$ is cyclic. According to Lemma 3.6, we have
${\mathbb{Q}}[G]^{\mathrm{rig}}\cong{\mathbb{Q}}(\zeta_{n}),$
where $n$ is the order of $G$; therefore, either $G$ is trivial, or $p$ divides $n$. It follows that, $E_{1}$ and $E_{2}$ are isogenous, and
$\operatorname{End}^{\circ}(A)\cong M(2,\operatorname{End}^{\circ}(E_{1})).$
According to Theorem 2.8, if $A$ is simple, then $\operatorname{End}^{\circ}(A)$ belongs to the following list:
  • ${\mathbb{H}}_{\infty}({\mathbb{Q}}(\sqrt{p}))$;
  • ${\mathbb{Q}}(\zeta_{8}\sqrt{p})$;
  • ${\mathbb{Q}}(\zeta_{6}\sqrt{p})$;
  • ${\mathbb{Q}}(\zeta_{12}\sqrt{p})$;
  • ${\mathbb{Q}}(\sqrt{5},\sqrt{-10-10\sqrt{5}})$, and $p=5$;
  • ${\mathbb{Q}}(\sqrt{3},\sqrt{-4-2\sqrt{3}})$, and $p=2$.
It is straightforward to check that in the last two cases, $\operatorname{End}^{\circ}(A)$ does not contain a cyclotomic field. We proved that the Weil polynomial $f=f_{A}$ of $A$ is one of the polynomials from cases $(1)-(6)$.
Let $A$ be an abelian variety with the Weil polynomial $f$. In the first two cases $L=\operatorname{End}^{\circ}(A)$ is a field:
  1. (1)
    $L={\mathbb{Q}}(\zeta_{8}\sqrt{p})={\mathbb{Q}}(\zeta_{4},\sqrt{2p})$, where $p\neq 2$;
  2. (2)
    $L={\mathbb{Q}}(\zeta_{6}\sqrt{p})={\mathbb{Q}}(\zeta_{3},\sqrt{p})$ or $L={\mathbb{Q}}(\zeta_{12}\sqrt{p})={\mathbb{Q}}(\zeta_{3},\sqrt{3p})$, where $p\neq 3$;
Clearly, $G$ is cyclic of order $4$ in case $(1)$, and of order $3$ or $6$ in case $(2)$. According to Corollary 3.9, in these cases there exists an abelian surface with a rigid action of $G$. By Lemmata 6.6, and 6.5, there exists an isogeny to an abelian surface with a rigid and symplectic action of $G$.
In case $(3)$ we have
${\mathbb{H}}=\operatorname{End}^{\circ}(A)\cong{\mathbb{H}}_{\infty}({\mathbb{Q}}(\sqrt{p})).$
If $G$ contains a cyclic subgroup of order $n\in\{5,8,12\}$, then there exists a homomorphism from ${\mathbb{Q}}(\zeta_{n})$ to ${\mathbb{H}}$. Dimension counting shows that the image contains ${\mathbb{Q}}(\sqrt{p})$, therefore $p$ divides $n$. A contradiction. According to Theorem 4.3, the group $G$ is either cyclic of order $3,4,6$, or isomorphic to $Q_{8}$, $Q_{12}$, or $\operatorname{SL}_{2}({\mathbb{F}}_{3})$.
We are going to construct rigid actions for these groups. According to Theorem 2.2, there are homomorphisms from ${\mathbb{Q}}[\zeta_{3}]$ and ${\mathbb{Q}}[\zeta_{4}]$ to ${\mathbb{H}}$. By Lemma 6.5, there exists an isogeny from $A$ to an abelian surface with a rigid and symplectic action of $G$. According to Corollary 3.9, there is an action of $Q_{8}$ or $\operatorname{SL}_{2}({\mathbb{F}}_{3})$ on some variety in the isogeny class of $A$ if and only if there is a homomorphism of ${\mathbb{H}}_{2}$ to ${\mathbb{H}}$. By Corollary 2.4, there is such a homomorphism if and only if $2$ does not split in ${\mathbb{Q}}(\sqrt{p})$. In the same way, one proves that $Q_{12}$ acts on some variety in the isogeny class of $A$ if and only if $3$ does not split in ${\mathbb{Q}}(\sqrt{p})$. According to Lemmata 6.7, and 6.8 these actions are symplectic.
In case $(4)$ the situation is similar:
${\mathbb{H}}=\operatorname{End}^{\circ}(A)\cong M_{2}({\mathbb{Q}}(\sqrt{-p})).$
In particular, the argument from the proof of case $(3)$ gives the same list of groups. According to Corollary 3.9 and Theorem 2.2, there are homomorphisms from ${\mathbb{Q}}[\zeta_{3}]$ and ${\mathbb{Q}}[\zeta_{4}]$ to ${\mathbb{H}}$. We use Corollary 3.9 and Corollary 2.4 again and find that there is an action of $Q_{8}$ or $\operatorname{SL}_{2}({\mathbb{F}}_{3})$ on some variety in the isogeny class of $A$ if and only if $2$ does not split in ${\mathbb{Q}}(\sqrt{-p})$, and $Q_{12}$ acts on some variety in the isogeny class of $A$ if and only if $3$ does not split in ${\mathbb{Q}}(\sqrt{-p})$.
Cases $(5)$ and $(6)$ can be treated in the same way. ∎
Theorem 7.11.
Assume that $k={\mathbb{F}}_{q}$, where $q=p^{2r+1}$, and $p>2$. There exists a Kummer surface $X$ over $k$ corresponding to an abelian surface $A$ with the following parameters.
$\operatorname{Tr}_{X}$ $Z_{X}(t)$ $G$ $p$ $f_{A}(t)$
$20$ $21^{21},$ 2 $Q_{8}$ $p\equiv 3(4)$ $(t^{2}-q)^{2}$
$18$ $1^{20},2^{2}$ $C_{4}$ $p\equiv 1(4)$ $(t^{2}-q)^{2}$
$18$ $1^{20},2^{2}$ $C_{2}$ $p\equiv 3(4)$ $(t^{2}+q)^{2}$
$14$ $1^{18},2^{4}$ $C_{2}$ $p\equiv 1(4)$ $(t^{2}-q)^{2}$
$10$ $1^{16},2^{6}$ $C_{2}$ $p\equiv 3(4)$ $(t^{2}+q)^{2}$
$8$ $1^{15},2^{7}$ $C_{4}$ $p\equiv 1(4)$ $(t^{2}-q)^{2}$
$6$ $1^{14},2^{8}$ $C_{2}$ $p>2$ $(t^{2}+q)^{2}$
$2$ $1^{12},2^{10}$ $C_{2}$ $p>2$ $(t^{2}-q)^{2}$
00 $1^{6},2^{4},3^{8},6^{4}$ $C_{2}$ $p>2$ $t^{4}-qt^{2}+q^{2}$
Proof.
If $G=C_{2}$, we apply [Ry14, Theorem 7.1]. If $p\equiv 1\bmod 4$, then, according to Proposition 7.3, there exists a supersingular surface $X$ with $\operatorname{Tr}_{X}=18$, and a supersingular surface with the trace equal to $8$.
Assume that $G=Q_{8}$, and $f_{A}(t)=(t^{2}-q)^{2}$. According to Theorem 7.10, there exists an abelian surface $A$ with the Weil polynomial $f$ and with a rigid and symplectic action of $Q_{8}$. By Lemma 7.2, the characteristic polynomial of Frobenius action on $H^{2}(\bar{A},{\mathbb{Q}}_{2})^{Q_{8}}$ is equal to $(t-q)^{2}(t+q)$. We construct an abelian surface $B$ with a $2$-isogeny $B\to A$ and an action of $Q_{8}$ such that each orbit of the action of Frobenius on $B[2]({\bar{k}})$ is a subset of an orbit of $Q_{8}$. According to Proposition 5.5 and Lemma 7.1, the zeta function of the Kummer surface $K(B,Q_{8})$ is $(1^{21},2)$.
The algebra $\operatorname{End}^{\circ}(A)$ is generated by ${\mathbb{H}}_{2}\cong{\mathbb{Q}}[Q_{8}]^{\mathrm{rig}}$ over its center ${\mathbb{Q}}(\sqrt{p})$. Let $i\in{\mathbb{H}}_{2}$ be the image of an element of $Q_{8}$ of order $4$, then there is a relation $(iF)^{2}=-q$. Let $T$ be a free submodule of $T_{2}(A)$ over the algebra $R={\mathbb{Z}}_{2}[i,y]$, where $y=(1+iF)/2$. Note that $T$ is $G$ invariant. According to Lemma 2.5, there exists $B$ such that $T_{2}(B)\cong T$. In particular, $F$ is equal to $i$ on $B[2]({\bar{k}})$. ∎